### v1

Abstract
In this paper we study a class of weighted estimands, which we define as parameters that can be expressed as weighted averages of the underlying heterogeneous treatment effects. The popular ordinary least squares (OLS), two-stage least squares (2SLS), and two-way fixed effects (TWFE) estimands are all special cases within our framework. Our focus is on answering two questions concerning weighted estimands. First, under what conditions can they be interpreted as the average treatment effect for some (possibly latent) subpopulation? Second, when these conditions are satisfied, what is the upper bound on the size of that subpopulation, either in absolute terms or relative to a target population of interest? We argue that this upper bound provides a valuable diagnostic for empirical research. When a given weighted estimand corresponds to the average treatment effect for a small subset of the population of interest, we say its internal validity is low. Our paper develops practical tools to quantify the internal validity of weighted estimands.
Keywords: internal validity, ordinary least squares, representativeness, treatment effects, two-stage least squares, two-way fixed effects, weakly causal estimands, weighted estimands
JEL classification: C20, C21, C23, C26
1 Introduction
Estimating average treatment effects is an important objective of empirical research in many areas of economics. Applied researchers usually believe that treatment effects are heterogeneous, which means that they vary between units with different observable and unobservable characteristics. Yet, many researchers also favor using well-established estimation methods that were not originally designed with treatment effect heterogeneity in mind. These methods may be chosen because of their computational simplicity, comparability across studies, effectiveness at incorporating high-dimensional covariates, and other reasons. In turn, these methods often lead to estimands that can be represented as weighted averages of the underlying treatment effects of interest.
For example, consider a scenario where unconfoundedness holds given covariates $X$. Let treatment $D$ be binary, $(Y(1),Y(0))$ be potential outcomes, and let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$ be the conditional average treatment effect, or CATE, for covariate value $X$. Following Angrist (1998), if we additionally assume that $\mathbb{E}[D\mid X]$ is linear in $X$, the regression of $Y$ on a constant, treatment $D$, and covariates $X$ yields a coefficient on $D$ that can be written as
$\beta_{\text{OLS}}=\frac{\mathbb{E}[\operatorname{var}(D\mid X)\tau_{0}(X)]}{\mathbb{E}[\operatorname{var}(D\mid X)]},$
a weighted average of CATEs with nonnegative weights that integrate to 1.
In this paper we are concerned with a general class of weighted estimands that can be expressed as follows:
$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$ (1.1)
where $W_{0}\in\{0,1\}$ is an indicator for a subpopulation, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ are the CATEs given covariates $X$ in subpopulation $W_{0}$, $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$ is the probability of being in subpopulation $W_{0}$ given $X$, and $a(X)$ is an identified weight function satisfying $\mathbb{E}[a(X)w_{0}(X)]>0$. The regression estimand above belongs to this class, which can be seen by letting $W_{0}=1$ with probability 1, and letting the weight function $a(X)$ be the conditional variance of treatment given covariates. Under some assumptions, this class also includes the two-stage least squares (2SLS) and two-way fixed effects (TWFE) estimands in instrumental variables and difference-in-differences settings, as well as many other parameters.
There are two main questions that this paper seeks to answer. The first is whether, and under what circumstances, the estimand in (1.1) corresponds to an average treatment effect of the form $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is an indicator for a (possibly latent) subpopulation of $W_{0}$. An affirmative answer to this question would endow a specific weighted estimand with some degree of validity as a causal parameter, given that it would then measure the average effect of treatment for a subset of all units.
The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter. To do this, we characterize the size, and the size relative to $W_{0}$, of subpopulations $W^{*}$ associated with the estimand in (1.1). More plainly, we ask how large $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can be in the representation $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units. If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target. If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$. Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function. If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form. We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs. In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function. While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be of interest—we consider this problem to be of independent theoretical interest. Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the impact of the nonuniform weights in $\mu(a,\tau_{0})$ in a given application.
In the first case, when the CATE function is unrestricted, we formally show that $\mu(a,\tau_{0})$ can be written as the average treatment effect for a subpopulation of $W_{0}$ if and only if $a(X)\geq 0$ with probability 1 given $W_{0}=1$. The contrapositive of this statement is that the incidence of “negative weights,” i.e. $\mathbb{P}(a(X)<0)>0$, implies that $\mu(a,\tau_{0})$ cannot be represented as an average treatment effect for some subpopulation uniformly in $\tau_{0}$. In a related contribution, Blandhol, Bonney, Mogstad, and Torgovitsky (2022) have shown that, for estimands that do not depend on potential outcome levels, the lack of negative weights is a sufficient and necessary condition for the weighted estimand to be “weakly causal,” that is, to guarantee that the sign of $\tau_{0}$ will be preserved whenever it is uniform across all units. We also provide simple expressions for the maxima of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. We propose an analog estimator for these objects and briefly discuss associated estimation and inferential issues.
In the second case, when the CATE function is assumed to be known, we show that $\mu(a,\tau_{0})$ can be written as an average treatment effect whenever it lies in the convex hull of CATE values, a weaker criterion than having nonnegative weights. The maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ now depend on $\tau_{0}$, and can be obtained via linear programming when $X$ is discrete. We show the solution to this linear program also admits a closed-form expression even when the support of $X$ includes discrete, continuous, and mixed components. This expression can be used to derive plug-in estimators.
Literature Review
This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models. Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Blandhol, Bonney, Mogstad, and Torgovitsky (2022), Słoczyński (2022), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
A common view in much of this literature, traceable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be positive. For example, Sun and Abraham (2021) explicitly associate “reasonable weights” with weights that “sum to one and are non-negative.” Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is necessary and sufficient for an estimand to be “weakly causal,” that is, to guarantee sign preservation when all treatment effects have the same sign. In this paper we focus on the related problem of whether a weighted estimand can be written as the average treatment effect for some (possibly latent) subpopulation. While the lack of negative weights is essential in our framework when the CATE function is unrestricted, negative weights and nonuniform weights play a similar role when the CATE function is assumed to be known, at least as long as the weighted estimand lies in the convex hull of CATE values. This point is related to the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
Some papers focus on weighted averages of heterogeneous treatment effects as legitimate targets in their own right rather than as probability limits of existing estimators. Hirano, Imbens, and Ridder (2003) introduce the class of weighted average treatment effects, which are a subclass of the more general class of estimands in (1.1). Li, Morgan, and Zaslavsky (2018) discuss the connection between weighted average treatment effects and implicit target subpopulations. However, the internal validity and representativeness of weighted estimands have received very little attention to date.
One exception is Aronow and Samii (2016), who explicitly acknowledge that the OLS estimand, like the local average treatment effect of Imbens and Angrist (1994), corresponds to the average effect for a “highly specific subpopulation” rather than the entire population, and consequently is not necessarily representative of that population. Then, Aronow and Samii (2016) focus on whether mean covariate values are similar in the entire sample and in the “effective sample” used by OLS. We focus on the size of the implicit subpopulation, which is different and complementary.
Another exception is Miller, Shenhav, and Grosz (2023), who focus on (one-way) fixed effects estimands and argue that it is problematic if “switchers,” that is, fixed-effect groups with nonzero variation in treatment, are a small subset of the sample. They also recommend that applied researchers report the sample size when limited to “switcher groups.” In this paper we build a general framework to study the internal validity and representativeness of weighted estimands, with the fixed effects estimand (equivalent to OLS) as a special case. We argue, similar to Miller, Shenhav, and Grosz (2023), that if a given weighted estimand corresponds to the average treatment effect for a small subpopulation, then it may not be an appropriate target parameter, unless that subpopulation is interesting in its own right.
Plan of the Paper
We organize the paper as follows. In Section 2, we discuss several examples of commonly used estimands that satisfy the representation in (1.1). In Section 3, we develop our theoretical framework and examine the conditions under which the estimand in (1.1) has a causal representation as an average treatment effect over a population. In Section 4, we establish our main results on the absolute and relative size of subpopulations associated with the estimand in (1.1), which we propose as measures of representativeness and internal validity of weighted estimands. We also generalize our results to subpopulations and revisit the examples from Section 2. In Section 5, we briefly discuss estimation and inference for the proposed measures. In Section 6, we provide an empirical application to the effects of unilateral divorce laws on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021). In Section 7, we conclude. Appendix contains our proofs as well as several additional results and derivations.
2 Applications
Here we consider three identification strategies where commonly used estimands follow the structure of equation (1.1).
2.1 Unconfoundedness
In this setup, we have a binary treatment $D\in\{0,1\}$, potential outcomes $(Y(1),Y(0))$, covariate vector $X$, and realized outcome $Y=Y(D)$. We make the following two assumptions.
Assumption 2.1 (Unconfoundedness).
Let
  1. 1.
    Conditional independence: $(Y(1),Y(0))\perp\!\!\!\perp D\mid X$;
  2. 2.
    Overlap: $p(X)\equiv\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1). The following proposition summarizes Angrist’s (1998) result.
Proposition 2.1.
Let Assumption 2.1 hold and let $p(X)$ be linear in $X$. Then
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[p(X)(1-p(X))\cdot\mathbb{E}[Y(1)-Y(0)\mid X]]}{\mathbb{E}[p(X)(1-p(X))]}.$
The linearity assumption can be removed if we instead regress $Y$ on $(1,D,h(X))$ where $h(X)$ is a vector of functions of $X$ such that $p(X)$ is in their linear span. The overlap assumption can also be weakened since it is not required for $\beta_{\text{OLS}}$ to be defined.
Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)]}{\mathbb{E}[a(X)]},$
where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. Here we implicitly set $W_{0}=1$ with probability 1. Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$. Note that $\beta_{\text{OLS}}=\text{ATE}\equiv\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case when $p(X)$ or $\tau_{0}(X)$ is constant.
An alternative representation of this estimand can be obtained by focusing on the subpopulation of treated units, $D=1$. Let $W_{0}=D$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,X]=\mathbb{E}[Y(1)-Y(0)\mid X]$, which follows from conditional independence, and let $\tilde{a}(X)=1-p(X)$. Then, we can write
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]}=\frac{\mathbb{E}[\tilde{a}(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[\tilde{a}(X)w_{0}(X)]},$
where $w_{0}(X)=\mathbb{P}(D=1\mid X)=p(X)$. Yet another representation can be obtained when focusing on the subpopulation of untreated units by letting $W_{0}=1-D$. We omit details for brevity.
2.2 Instrumental Variables
Consider an endogenous binary treatment $D\in\{0,1\}$ and a binary instrument $Z\in\{0,1\}$. Potential treatments, denoted by $(D(1),D(0))$, are linked to the realized treatment through $Z$, that is, $D=D(Z)$. Potential outcomes, $Y(d,z)$ for $d,z\in\{0,1\}$, may depend on both $D$ and $Z$ in the absence of an exclusion restriction. Let $Y=Y(D,Z)$ be the realized outcome. As before, let $X$ denote covariates. We make the following assumptions.
Assumption 2.2 (Instrument validity).
We have
  1. 1.
    Exogeneity: $(Y(0,0),Y(1,0),Y(0,1),Y(1,1),D(1),D(0))\perp\!\!\!\perp Z\mid X$;
  2. 2.
    Exclusion: $\mathbb{P}(Y(d,0)=Y(d,1))=1$ for $d\in\{0,1\}$;
  3. 3.
    First stage: $e(X)\equiv\mathbb{P}(Z=1\mid X)\in(0,1)$ and $\mathbb{P}(D(1)=1\mid X)\neq\mathbb{P}(D(0)=1\mid X)$ almost surely.
We also make one of the following two nested monotonicity assumptions.
Assumption 2.3 (Strong monotonicity).
$\mathbb{P}(D(1)\geq D(0)\mid X)=1$ almost surely.
Assumption 2.4 (Weak monotonicity).
There exists a subset of the support of $X$ such that $\mathbb{P}(D(1)\geq D(0)\mid X)=1$ on it and $\mathbb{P}(D(1)\leq D(0)\mid X)=1$ on its complement.
The first instrumental variables estimand we consider was originally studied by Angrist and Imbens (1995). In addition to the assumptions on the instrument, suppose that the model for $X$ is saturated, with $K$ possible combinations of covariate values, i.e. let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, where $\text{supp}(\cdot)$ is used to denote the support. Let $X_{S}=\left(1,\mathbbm{1}(X=x_{1}),\ldots,\mathbbm{1}(X=x_{K-1})\right)$ and $Z_{S}=\left(Z,Z\mathbbm{1}(X=x_{1}),\ldots,Z\mathbbm{1}(X=x_{K-1})\right)=ZX_{S}$, where $Z_{S}$ is the constructed instrument vector. The estimand in Angrist and Imbens (1995) is the following 2SLS estimand:
$\beta_{\text{2SLS}}\equiv\left[\left(\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}W_{S}\right]\right)^{-1}\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}Y\right]\right]_{1},$
where $W_{S}=\left(D,X_{S}\right)$, $Q_{S}=\left(Z_{S},X_{S}\right)$, and $\left[\cdot\right]_{k}$ denotes the $k$th element of the corresponding vector. This estimand has been studied by Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022), and the specific representation in Proposition 2.2 follows Słoczyński (2020).
Proposition 2.2.
Let Assumptions 2.2 and 2.4 hold, and let $X$ be discrete with finite support. Then
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}\left[e(X)(1-e(X))\cdot\mathbb{P}\left(D(1)\neq D(0)\mid X\right)^{2}\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]\right]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}.$
This means that we can write $\beta_{\text{2SLS}}$ as
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
where $W_{0}=\mathbbm{1}(D(1)\neq D(0))$ is the population of compliers/defiers, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and $a(X)=e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, a nonnegative weight function. Note that $\beta_{\text{2SLS}}=\text{LATE}\equiv\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0)]$ if and only if $a(X)$ is uncorrelated with $\tau_{0}(X)$ given $D(1)\neq D(0)$.
The practical limitation of focusing on $\beta_{\text{2SLS}}$ is that applied researchers rarely create additional instruments by interacting the original instrument with covariates (cf. Blandhol, Bonney, Mogstad, and Torgovitsky, 2022), which is how $Z_{S}$ is constructed to obtain $\beta_{\text{2SLS}}$ above. A more practically relevant estimand is the “just-identified” IV estimand,
$\beta_{\text{IV}}\equiv\left[\left(\mathbb{E}\left[Q^{\prime}W\right]\right)^{-1}\mathbb{E}\left[Q^{\prime}Y\right]\right]_{1},$
where $Q=\left(Z,X\right)$ and $W=\left(D,X\right)$. To introduce one of the representations of $\beta_{\text{IV}}$ below, define
$c(X)\equiv\operatorname{sign}\Big{(}\mathbb{P}\left[D(1)\geq D(0)\mid X\right]-\mathbb{P}\left[D(1)\leq D(0)\mid X\right]\Big{)},$
where $\operatorname{sign}(\cdot)$ is the sign function. We also make the following “rich covariates” assumption on the instrument propensity score, which is implied by the saturated specification in Proposition 2.2.
Assumption 2.5 (Rich covariates).
$e(X)$ is linear in $X$.
Under the instrument validity assumption, the rich covariates assumption, and either monotonicity assumption, Słoczyński (2020) obtains the following representations for the “just-identified” IV estimand.
Proposition 2.3.
Let Assumptions 2.2, 2.4, and 2.5 hold. Then
$\beta_{\text{IV}}=\frac{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]]}{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)]}.$
If Assumptions 2.2, 2.3, and 2.5 hold instead, then
$\beta_{\text{IV}}=\frac{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)]}.$
It follows that we can write $\beta_{\text{IV}}$ as
$\beta_{\text{IV}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}$
under either monotonicity assumption. Under weak monotonicity, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and possibly negative weights $a(X)=c(X)\cdot e(X)(1-e(X))$. Under strong monotonicity, $W_{0}=\mathbbm{1}(D(1)>D(0))$, $w_{0}(X)=\mathbb{P}(D(1)>D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and nonnegative weights $a(X)=e(X)(1-e(X))\geq 0$.
2.3 Difference-in-Differences
Now suppose units are observed for $T$ periods and, for $t\in\{1,\ldots,T\}$, denote binary treatment $D_{t}\in\{0,1\}$, potential outcomes $(Y_{t}(1),Y_{t}(0))$, and realized outcome $Y_{t}=Y_{t}(D_{t})$. We assume units are untreated prior to period $G\in\mathcal{G}=\{2,3,\ldots,T\}\cup\{+\infty\}$, receive the treatment in period $G$, and remain treated thereafter. We assume no units are treated in the first time period. This may include a group that remains untreated throughout, for which $G=+\infty$. Thus, $D_{t}=\mathbbm{1}(G\leq t)$. The panel is balanced, that is, no group appears or disappears over time.
The two-way fixed effects estimand is often used in this setting, and consists of regressing the outcome on the treatment indicator, group indicators, and period indicators. Using partitioned regression results, it is defined as
$\beta_{\text{TWFE}}\equiv\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}Y_{t}\right]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}^{2}\right]},$
where $\ddot{D}_{t}=D_{t}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]$.
We assume a version of parallel trends most similar to the one in de Chaisemartin and D’Haultfœuille (2020).
Assumption 2.6 (Difference-in-differences).
We have
  1. 1.
    $\mathcal{G}\equiv\text{supp}(G)=\{2,3,\ldots,T\}\cup\{+\infty\}$;
  2. 2.
    For all $t\in\{2,\ldots,T\}$ and $g,g^{\prime}\in\mathcal{G}$, we have that $\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g]=\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g^{\prime}]$.
We use a proposition that is essentially a special case of Theorem 1 in de Chaisemartin and D’Haultfœuille (2020) to obtain a representation of the two-way fixed effects estimand as a weighted average.
Proposition 2.4.
Let Assumption 2.6 hold. Then
$\beta_{\text{TWFE}}=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\cdot\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid G,D_{t}=1]\Big{]}}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\Big{]}}.$
We show the above representation satisfies equation (1.1) by introducing an auxiliary variable $P$ that is uniformly distributed on $\{1,\ldots,T\}$ independently from $\{(Y_{t}(0),Y_{t}(1),G)\}_{t=1}^{T}$. This period variable denotes the time period and we use it to define $(Y(1),Y(0))\equiv(Y_{P}(1),Y_{P}(0))$, $Y=Y_{P}$, and $D=D_{P}$, which are potential outcomes, the realized outcome, and treatment at a random time period, respectively.
Letting $X=(G,P)$, this means we can write $\beta_{\text{TWFE}}$ as
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
where $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G,P)\in\{0,1\}$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G,P]$, and the weight function $a(X)=1-\mathbb{P}(D=1\mid G)-\mathbb{P}(D=1\mid P)+\mathbb{P}(D=1)$ is not generally nonnegative. [^2] Here, $\tau_{0}(X)$ is what Callaway and Sant’Anna (2021) call “the group-time average treatment effect.” A nonnegative weight function can be obtained under the assumption that group-level average treatment effects are constant over time. This property was described in Goodman-Bacon (2021, Section 3.1.1) and a representation of the two-way fixed effects estimand is given in his equation (16). The following proposition yields a simpler (but equivalent) expression for the weights in our setting.
Proposition 2.5.
Let Assumption 2.6 hold and let $\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid D=1,G]=\mathbb{E}[Y_{s}(1)-Y_{s}(0)\mid D=1,G]$ for any $s,t\in\{1,\ldots,T\}$. Then
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid D=1,G]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]},$
where $a_{B}(g)=\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\geq 0$ for $g\in\{2,\ldots,T\}$.
As is the case of the representation in Proposition 2.4, the two-way fixed effects estimand in Proposition 2.5 satisfies the representation in (1.1), with $X=G$, $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G]$, and the weight function $a(X)=a_{B}(G)\geq 0$. This weight function is derived in the proof of the proposition, and we show it is equivalent to equation (16) in Goodman-Bacon (2021) in Appendix D.
3 Causal Representation of Weighted Estimands
We now consider a class of weighted estimands that subsumes all our previous examples. In this section, we show necessary and sufficient conditions for an estimand of this class to have a causal representation as an average treatment effect over a population. Then, we compute the largest size of a population that is compatible with the weighted estimand. The size of this population relative to $\mathbb{P}(W_{0}=1)$ will be a measure of the internal validity of our estimand relative to the target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. The absolute size of this population will be a measure of the representativeness of the estimand with respect to the population distribution the data is drawn from.
Recall the earlier setting where we let $D\in\{0,1\}$ denote a binary treatment variable, and let $(Y(1),Y(0))$ denote the corresponding potential outcomes under treatment and control, respectively. Let $X\in\text{supp}(X)\subseteq\mathbb{R}^{d_{X}}$ denote a $d_{X}$-vector of covariates. We suppose that $(Y(1),Y(0),D,X)$ are drawn from a common population distribution $F_{Y(1),Y(0),D,X}$.
Let $W_{0}\in\{0,1\}$ be an indicator variable used to denote a subpopulation $\{W_{0}=1\}$. Throughout this paper, we assume that $\mathbb{P}(W_{0}=1)>0$, so that this subpopulation has a positive mass, which avoids technical issues associated with conditioning on zero-probability events. For example, this subpopulation can be the entire population by setting $W_{0}=1$ almost surely, the treated units by setting $W_{0}=D$, or units complying with binary instrument $Z$ by setting $W_{0}=\mathbbm{1}(D(1)>D(0))$. Let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ denote the conditional average treatment effect given $X$ in subpopulation $W_{0}$. This is defined for all values of $X$ such that $w_{0}(X)>0$.
Also recall the weighted estimands of equation (1.1):
$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$
where $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$. [^3] While $\tau_{0}(X)$ is only defined for $w_{0}(X)>0$, we set $\tau_{0}(X)w_{0}(X)=0$ when $w_{0}(X)=0$.
The estimands we consider have the above representation and satisfy the following regularity conditions.
Assumption 3.1 (Regularity).
Let $\mathbb{E}[\tau_{0}(X)^{2}]<\infty$, $\mathbb{E}[a(X)^{2}]<\infty$, and $\mathbb{E}[a(X)\mid W_{0}=1]>0$.
The first two restrictions are weak regularity assumptions that ensure the finiteness of the numerator of $\mu(a,\tau_{0})$. We rule out $\mathbb{E}[a(X)\mid W_{0}=1]=0$ since it implies the estimand does not exist. The estimand in (1.1) is unchanged if the sign of $a(X)$ is reversed, so $\mathbb{E}[a(X)\mid W_{0}=1]>0$ is a sign normalization.
3.1 Alternative Representations of Weighted Estimands
The weighted estimands of equation (1.1) can also be written as a weighted sum when $X$ is discrete, or an integral when $X$ is continuous. In the discrete case, let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, $W_{0}=1$ almost surely, and let $p_{k}\equiv\mathbb{P}(X=x_{k})>0$ for $k=1,\ldots,K$. Then,
$\mu(a,\tau_{0})$ $=\sum_{k=1}^{K}\omega_{k}\tau_{0}(x_{k})\qquad\text{ where }\qquad\omega_{k}=\frac{a(x_{k})p_{k}}{\sum_{l=1}^{K}a(x_{l})p_{l}},$ (3.1)
which are weights that sum to one. The representations in (1.1) and (3.1) are equivalent as we can obtain $a(x_{k})$ (up to scale) as the ratio $\omega_{k}/p_{k}$, and $\omega_{k}$ is defined as a function of $(a(\cdot),p_{\cdot})$ in equation (3.1) above.
From (3.1), we can see that $a(x_{k})$ being constant ensures $\omega_{k}=p_{k}$, or that the estimand is the ATE. Moreover,
$\frac{a(x_{k})}{a(x_{k^{\prime}})}=\frac{\omega_{k}}{\omega_{k^{\prime}}}\big{/}\frac{p_{k}}{p_{k^{\prime}}},$
which is the ratio of the relative weights of covariate cells $\{X=x_{k}\}$ and $\{X=x_{k^{\prime}}\}$ in the estimand $(\omega_{k}/\omega_{k^{\prime}})$ and in the population $(p_{k}/p_{k^{\prime}})$. The inequality $a(x_{k})>a(x_{k^{\prime}})$ indicates that covariate cell $\{X=x_{k}\}$ is overweighted relative to $\{X=x_{k^{\prime}}\}$, when compared to their relative weights in the population.
Alternatively, consider the case where $X$ is continuously distributed with density [^4] With respect to the Lebesgue measure. $f_{X}$. Still assuming $W_{0}=1$ almost surely, we can write
[…]
is a weight function that integrates to 1. We focus on the representation in equation (1.1) since it seamlessly accommodates discrete, continuous, and mixed covariates.
3.2 Regular Subpopulations
The first question we address is whether an estimand defined by (1.1) can be represented as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is binary and $\{W^{*}=1\}$ characterizes a subpopulation of $\{W_{0}=1\}$. Formally, $\{W^{*}=1\}$ forms a subpopulation if $\{W^{*}=1\}\subseteq\{W_{0}=1\}$ or, equivalently, if $W^{*}\leq W_{0}$ almost surely.
We impose some structure on this problem by restricting how these subpopulations may be formed. We will consider what we call “regular subpopulations,” which we define here.
Definition 3.1.
Let $W^{*}\in\{0,1\}$ such that $\mathbb{P}(W^{*}=1)>0$. Say $\{W^{*}=1\}$ is a regular subpopulation of $\{W_{0}=1\}$ if
  1. 1.
    (Inclusion) $\mathbb{P}(W^{*}=1\mid W_{0}=0)=0$;
  2. 2.
    (Conditional Independence) $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$.
For convenience, we will abbreviate this as “$W^{*}$ is a regular subpopulation of $W_{0}$”. We denote the set of regular subpopulations of $W_{0}$ as
$\text{SP}(W_{0})$ =$=\{W^{*}\in\{0,1\}:W^{*}\text{ is a regular subpopulation of }W_{0}\}.$
We consider subpopulations with positive masses that are subsets of $\{W_{0}=1\}$. The second and main requirement is that regular subpopulations do not depend on potential outcomes when conditioning on $X$ and the original population $W_{0}=1$. While this may seem restrictive, it allows for rich and natural classes of subpopulations. For example, consider the unconfoundedness restriction of Section 2.1 and let $W_{0}$ be the entire population, i.e. $\mathbb{P}(W_{0}=1)=1$. In this case, regular subpopulations must satisfy $\;\;W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X$, or be unconfounded. Regular subpopulations include the population of all treated (or untreated) individuals, i.e. $W^{*}=D$ (or $W^{*}=1-D$), and any subpopulation characterized by a subset of $\text{supp}(X)$. More generally, it includes any subpopulation that can be described through a combination of $(D,X,U)$ where $U$ is independent from $(Y(1),Y(0),X)$. For example, a subpopulation characterized by “fraction $a(x)$ of units with covariate $X=x$ for all $x\in\text{supp}(X)$” can be constructed as $W^{*}=\mathbbm{1}(U\leq a(X))$ where $U\sim\text{Unif}(0,1)$ is independent from $(Y(1),Y(0),X)$.
The conditional independence requirement rules out subpopulations that directly depend on the potential outcomes such as $W^{*}=\mathbbm{1}(Y(1)\geq Y(0))$, the subpopulation of those who benefit from treatment. Note that $\mathbb{P}(W^{*}=1\mid X)=\mathbb{P}(Y(1)\geq Y(0)\mid X)$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(Y(1)\geq Y(0))$ are not point-identified under unconfoundedness.
These particular subpopulations enjoy a number of useful properties. Two of them are characterized in the following proposition.
Proposition 3.1 (Properties of regular subpopulations).
Let $\mathbb{P}(W_{0}=1)>0$ and $W^{*}\in\text{SP}(W_{0})$.
  1. 1.
    Let $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)>0$. Then,
    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X].$ (3.3)
  2. 2.
    Let $\mathbb{E}[\tau_{0}(X)^{2}]<\infty$. Then,
    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0}),$ (3.4)
    where $\underline{w}^{*}(x)=\mathbb{P}(W^{*}=1\mid X=x,W_{0}=1)$.
The first part of this proposition shows that average effects within the original population $W_{0}$ and regular subpopulation $W^{*}$ are the same when conditioning on $X$. For example, this holds under unconfoundedness for the subpopulation of treated individuals, $W^{*}=D$. The second property allows us to write the average effect for $W^{*}=1$ using the same functional $\mu(\cdot,\cdot)$ that was used to characterize the class of estimands we analyze. This property will be used when studying the mapping between weighted estimands and average effects for regular subpopulations of $W_{0}$.
We conclude this subsection by showing that regular subpopulations are transitive, or that regular subpopulations of a regular subpopulation of $W_{0}$ are also regular subpopulations of $W_{0}$.
Lemma 3.1 (Transitivity of regular subpopulations).
Let $W^{*}$ be a regular subpopulation of $W^{\prime}$ and let $W^{\prime}$ be a regular subpopulation of $W_{0}$. Then, $W^{*}$ is a regular subpopulation of $W_{0}$.
3.3 Existence of a Causal Representation for Weighted Estimands
We now consider necessary and sufficient conditions for the weighted estimand $\mu(a,\tau_{0})$ to be written as the average treatment effect within a regular subpopulation of $W_{0}$. As we will show, these conditions depend on what is assumed about the function $\tau_{0}=\mathbb{E}[Y(1)-Y(0)\mid X=\cdot,W_{0}=1]$.
For example, if $\tau_{0}$ is constant in $X$, then any weighted estimand satisfying (1.1) equals $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average treatment effect within population $\{W_{0}=1\}$. This is the case even when the sign of weight function $a(X)$ varies with $X$. However, if $\tau_{0}$ is nonconstant, the existence of causal representations will depend on the weight function $a(X)$. Among other cases, we will consider the case where no restrictions are placed on function $\tau_{0}$, and in this case, the existence of a causal representation of $\mu(a,\tau_{0})$ will require the sign of $a(X)$ to be constant.
To formalize this, let $\mathcal{T}$ denote a class of functions such that $\tau_{0}\in\mathcal{T}$ and define
$\mathcal{W}(a;W_{0},\mathcal{T})$ =$=\{W^{*}\in\text{SP}(W_{0}):\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\text{ for all }\tau_{0}\in\mathcal{T}\}.$
This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all $\tau_{0}$ functions in the set $\mathcal{T}$. If it is empty, then estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$. We use this set to formally define a notion of uniform causal representation.
Definition 3.2.
A weighted estimand $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}$ if
$\mathcal{W}(a;W_{0},\mathcal{T})\neq\emptyset.$
Recall that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$, so $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$ if $\mu(a,\tau_{0})=\mu(\underline{w}^{*},\tau_{0})$ for all $\tau_{0}\in\mathcal{T}$. We examine further two main cases for the set $\mathcal{T}$.
3.3.1 Existence Uniformly in $\tau_{0}$
We begin by considering the largest class of functions in which $\tau_{0}$ lies: the class of all functions, subject to the moment condition in Assumption 3.1 that ensures the existence of $\mu(a,\tau_{0})$. We denote this class by
$\mathcal{T}_{\text{all}}$$\equiv\{\tau_{0}:\mathbb{E}[\tau_{0}(X)^{2}]<\infty\}.$
In this function class, the existence of a causal representation is equivalent to the weights being nonnegative.
Theorem 3.1.
Let Assumption 3.1 hold and let $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$ if and only if $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
A uniform (in $\mathcal{T}_{\text{all}}$) causal representation exists if and only if $a(X)$ is nonnegative on the support of $X\mid W_{0}=1$. To give some intuition on why the sign of $a(X)$ must be nonnegative with probability 1, partition $\text{supp}(X\mid W_{0}=1)$ in $\mathcal{X}^{+}\equiv\{x\in\text{supp}(X\mid W_{0}=1):a(x)\geq 0\}$ and $\mathcal{X}^{-}\equiv\{x\in\text{supp}(X\mid W_{0}=1):a(x)<0\}$. If $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$, then $\mathbb{P}(X\in\mathcal{X}^{+}\mid W_{0}=1)>0$ and $\mathbb{P}(X\in\mathcal{X}^{-}\mid W_{0}=1)>0$ under Assumption 3.1. Then, setting $\tau^{-}(X)=\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau^{+}(X)=\mathbbm{1}(X\in\mathcal{X}^{+})$ yields estimands of different signs, i.e. $\mu(a,\tau^{-})<0<\mu(a,\tau^{+})$, despite $\tau^{-}$ and $\tau^{+}$ being of the same sign. In other words, as is well known, nonnegative CATEs can be combined into a negative estimand when weights are negative. Hence, such an estimand is not weakly causal (Blandhol, Bonney, Mogstad, and Torgovitsky, 2022).
Conversely, if $a(X)\geq 0$, our proof constructively defines a regular subpopulation $W^{*}$ for which the average effect is equal to the weighted estimand $\mu(a,\tau_{0})$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. Let
$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0},$
where $U\sim\text{Unif}(0,1)\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$. This is a regular subpopulation of $W_{0}$ for which the probability of inclusion, conditional on $X$ and $W_{0}=1$, is proportional to $a(X)$. The condition $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$ restricts our attention to subpopulations with positive mass. From this construction, we can see that $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is proportional to $a(X)$, and therefore $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})=\mu(a,\tau_{0})$ uniformly in $\tau_{0}$.
3.3.2 Existence for a Given $\tau_{0}$
We now provide an existence result that requires the causal representation to exist only for the given $\tau_{0}$, rather than uniformly for $\tau_{0}$ in the larger set $\mathcal{T}_{\text{all}}$. This result’s assumptions depend on the nature of the CATE function $\tau_{0}$ in the population, whereas Theorem 3.1’s condition depended only on the weight function $a(X)$, and the nature of the covariates’ support. Thus, the distribution of the potential outcomes will have an impact on the existence of a causal representation given $\tau_{0}$. Using the notation from Definition 3.2, a causal representation exists if and only if $\mathcal{W}(a;W_{0},\{\tau_{0}\})\neq\emptyset$. The following theorem characterizes this existence.
Theorem 3.2.
Let Assumption 3.1 hold. Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if and only if
$\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})\equiv\{t\in\mathbb{R}:\mathbb{P}(\tau_{0}(X)\leq t\mid W_{0}=1)>0\text{ and }\mathbb{P}(\tau_{0}(X)\geq t\mid W_{0}=1)>0\}.$ (3.5)
The existence condition in Theorem 3.2 is weaker than the one in Theorem 3.1 since we no longer require this representation to be valid for all $\tau_{0}\in\mathcal{T}_{\text{all}}$, but rather for just one its element. The necessary and sufficient condition in this theorem is rather weak, since it only requires that the estimand is in the convex hull of the support of the CATEs. This means $\mu(a,\tau_{0})$ has a causal representation even with negative weights, as long as there are CATEs smaller and greater than $\mu(a,\tau_{0})$. We can see this support condition holds for all $\tau_{0}\in\mathcal{T}_{\text{all}}$ if and only if $\mu(a,\tau_{0})$ is in the support of $\tau_{0}(X)$ for all $\tau_{0}$. This is precisely the case when the weights $a(X)$ are nonnegative since it guarantees $\inf(\text{supp}(\tau_{0}(X)\mid W_{0}=1))\leq\mu(a,\tau_{0})\leq\sup(\text{supp}(\tau_{0}(X)\mid W_{0}=1))$ for any function $\tau_{0}$.
3.3.3 Intermediate Cases
Analyzing the representativeness of an estimand under no restrictions on $\tau_{0}$ could be viewed as unnecessarily conservative in some settings. At the other extreme, assuming knowledge of $\tau_{0}$ may be unrealistic, especially in scenarios where $X$ has many components which makes the estimation of $\tau_{0}$ more challenging. For example, some shape constraints may be known to hold for $\tau_{0}$. In some economic applications one may posit that $\tau_{0}$ is monotonic or convex in some components of $X$, or positive/negative over a subset of $\text{supp}(X\mid W_{0}=1)$. In these cases, the existence of a causal representation may occur under weaker conditions than Theorem 3.1, but stronger than Theorem 3.2. In particular, one may be able to relax the requirement that $a(X)\geq 0$ without requiring that $\tau_{0}$ be completely known to the researcher. The following proposition shows this is the case when $\tau_{0}(X)$ is assumed to be linear in $X$.
Proposition 3.2.
Let Assumption 3.1 hold and define
$\mathcal{T}_{\text{lin}}$ =$=\{\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$ if and only if
$\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}\in\text{conv}(\text{supp}(X\mid W_{0}=1)),$
where $\text{conv}(\cdot)$ denotes the convex hull.
The above proposition shows that once the class of CATE functions is restricted, the requirement that $a(X)$ be nonnegative is no longer needed for the existence of a uniform causal representation for an estimand. In particular, the requirement here is that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$ lies in the convex hull of the support of $X$ given $W_{0}=1$. When $X$ is scalar, this consists of an interval. This condition does not require $a(X)$ be nonnegative: for example, if $\text{supp}(X)=\{0,1,2\}$, $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$, and $W_{0}=1$ almost surely, then any combination of $(a(0),a(1),a(2))$ values such that $\frac{a(1)+2a(2)}{a(0)+a(1)+a(2)}\in[0,2]$ implies a causal representation. For example, $(a(0),a(1),a(2))=(1,-1,1)$ contains a negative element, and thus $\mathbb{P}(a(X)\geq 0)=2/3<1$, but $\frac{a(1)+2a(2)}{a(0)+a(1)+a(2)}=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation.
We consider another class of CATE functions that restricts their heterogeneity. For $K\geq 0$, let
$\mathcal{T}_{\text{BV}}(K)=\left\{\tau\in\mathcal{T}_{\text{all}}:\sup_{x,x^{\prime}\in\text{supp}(X\mid W_{0}=1)}|\tau(x)-\tau(x^{\prime})|\leq K\right\}.$
This function class bounds the variation of the CATE function. When $K=0$, the CATE function is constant, and thus equal to $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. When $K>0$, CATEs may differ in value, but the maximum discrepancy between two CATEs is bounded above by $K$. We show that restricting the CATEs to satisfy this bounded variation assumption does not remove the requirement that $a(X)$ be nonnegative, unless $K=0$, in which case all $a$ functions yield a causal representation uniformly in $\mathcal{T}_{\text{BV}}(0)$. We formalize this in the next proposition.
Proposition 3.3.
Let Assumption 3.1 hold. If $K>0$, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ if and only if $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$. If $K=0$, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(0))$ for all $a$.
The last two propositions show that the impact of restrictions on $\tau_{0}$ on the requirement that $a(X)$ be nonnegative depends on the nature of these restrictions. Generalizations to additional or empirically motivated function classes is left for future work.
4 Quantifying the Internal Validity of Weighted Estimands
Many estimands will admit causal representations, but their associated subpopulations $\{W^{*}=1\}$ will generally differ. Also, a weighted estimand may not always correspond to the target estimand a researcher is interested in, such as $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ if the average effect in $\{W_{0}=1\}$ is the parameter of interest. However, the set of $W^{*}$ corresponding to a weighted estimand can be used to understand how representative the weighted estimand is of the target. For example, we may seek estimands for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ attains values closest to 1, since they have a higher degree of internal validity with respect to the target $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. At one extreme, an estimand for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=1$ is perfectly representative of $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ and would be deemed to have the highest degree of internal validity for this target parameter.
Formally, we ask what is the sharp upper bound on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ for any regular subpopulation $W^{*}$ of $W_{0}$ such that the weighted estimand $\mu(a,\tau_{0})$ has a causal representation as the average treatment effect over subpopulation $W^{*}$. We denote this object using the following notation:
$\overline{P}(a,W_{0};\mathcal{T})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})}\mathbb{P}(W^{*}=1\mid W_{0}=1).$ (4.1)
We also set $\overline{P}(a,W_{0};\mathcal{T})=0$ when $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty. This object depends on the chosen function class $\mathcal{T}$, as did Theorems 3.1 and 3.2 in the previous section. Given the above terminology, we call $\overline{P}(a,W_{0};\mathcal{T})$ a measure of the internal validity of estimand $\mu(a,\tau_{0})$, assuming that $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target.
We can also compute the maximum value of $\mathbb{P}(W^{*}=1)$ across $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$, which equals $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)\cdot\mathbb{P}(W_{0}=1)$ since $W^{*}$ is a subpopulation of $W_{0}$. This maximum value of $\mathbb{P}(W^{*}=1)$ gives the internal validity of the weighted estimand with respect to target estimand $\mathbb{E}[Y(1)-Y(0)]$, the average treatment effect in the population from which the sample is drawn. It can also be understood as a measure of representativeness of the weighted estimand with respect to that population.
As earlier, we break down our results in two cases, the first being when $\tau_{0}$ is unrestricted.
4.1 Quantifying Internal Validity Uniformly in $\tau_{0}$
Without imposing any restrictions on the CATE function, except for the existence of second moments, the maximum value that $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can achieve is given by the following theorem.
Theorem 4.1.
Let Assumption 3.1 hold and let $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. Then
$\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})}\mathbb{P}(W^{*}=1\mid W_{0}=1)=\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\begin{cases}\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\sup(\text{supp}(a(X)\mid W_{0}=1))}&\text{ if }\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1\\ 0&\text{ if }\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1.\end{cases}$ (4.2)
Here we see that the maximum size of a subpopulation characterizing the estimand $\mu(a,\tau_{0})$ depends on $a(X)$ through two terms: the denominator of the estimand, $\mathbb{E}[a(X)\mid W_{0}=1]$, and the maximum value of $a(x)$ over the support of $X\mid W_{0}=1$. Note that $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is invariant to the scale of $a(X)$, as is the estimand $\mu(a,\tau_{0})$. To understand the supremum’s role in this expression, let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ and note that $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ is equivalent to writing
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0})$ (4.3)
for all $\tau_{0}$. Equation (4.3) holding for all $\tau_{0}$ requires $\underline{w}^{*}(X)$ be exactly proportional to $a(X)$. While $a(X)$’s range is unconstrained, $\underline{w}^{*}(X)$ must lie in $[0,1]$ to be a valid conditional probability. Since we seek to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we let $\underline{w}^{*}(X)$ be the largest multiple of $a(X)$ that lies in $[0,1]$ with probability 1, which is defined below:
$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0}\qquad\text{ and }\qquad\underline{w}^{*}(X)=\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}.$
Here, $U\sim\text{Unif}(0,1)$ and $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$. This population places relatively more weight on units with larger values of $a(X)$. Specifically, the population $\{W^{*}=1\}$ contains a random subset of $\{W_{0}=1\}$ where the probability of inclusion is proportional to $a(X)$. Thus, units with larger values of $a(X)$ are more likely to be included in $W^{*}$. All units in $\{W_{0}=1\}$ with $X$ such that $a(X)=\sup(\text{supp}(a(X)\mid W_{0}=1))$ are included in $W^{*}$, whereas no units where $a(X)=0$ are included.
Given a data-generating process, this bound can be computed without any knowledge of the outcome’s conditional distribution and depends solely on weight function $a(\cdot)$ and the distribution of $X\mid W_{0}=1$.
Remark 4.1 (Defining the target estimand from the weighted estimand).
Suppose we consider $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ to be the target estimand, where $W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0}$ is the subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ uniformly in $\tau_{0}$. For example, under Example 2.1, this consists of a subpopulation where the probability of inclusion is proportional to $\operatorname{var}(D\mid X)$, the conditional variance of treatment. If this subpopulation is the target, it would be reasonable to infer that the measure of internal validity of the estimand $\mu(a,\tau_{0})$ is the maximum value of 1. Indeed, this is the case because the estimand can be written as
$\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
and $\{W^{*}=1\}$ is trivially the largest regular subpopulation of $\{W^{*}=1\}$. This illustrates how the internal validity of an estimand is entirely dependent on the target estimand. On the other hand, the representativeness of the weighted estimand, as measured by the largest value of $\mathbb{P}(W^{*}=1)$, is less than one unless $W^{*}=W_{0}$ almost surely, or that the weighted estimand actually equals the ATE. Theorem 4.4 below can also be applied to obtain this intuition.
We now consider a simple example to give further intuition for Theorem 4.1.
4.1.1 Illustrative Example: A Single Binary Covariate
Consider an estimand $\mu(a,\tau_{0})$ where $W_{0}=1$ almost surely, $a(X)\geq 0$, and where $X$ is binary with support $\text{supp}(X)=\{0,1\}$. Let $p_{x}=\mathbb{P}(X=x)>0$ for $x\in\{0,1\}$. As in Section 3.1, the weighted estimand can be written as a linear combination of the two CATEs:
$\mu(a,\tau_{0})$ $=\frac{a(0)p_{0}}{\mathbb{E}[a(X)]}\tau_{0}(0)+\frac{a(1)p_{1}}{\mathbb{E}[a(X)]}\tau_{0}(1)\equiv\omega_{0}\tau_{0}(0)+\omega_{1}\tau_{0}(1).$
Let $\mathbb{E}[Y(1)-Y(0)]=\text{ATE}$ be the target estimand, which can be written as
ATE $=p_{0}\tau_{0}(0)+p_{1}\tau_{0}(1).$
If $a(0)=a(1)$, the relative weights placed on $\{X=0\}$ and $\{X=1\}$ by the estimand are equal to $p_{0}/p_{1}$, the ratio of the weights placed by the ATE. Therefore, the estimand equals the ATE and thus clearly has the maximum degree of internal validity with respect to the ATE. Applying Theorem 4.1, we can directly see that
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)]}{\sup_{x\in\{0,1\}}a(x)}=\frac{a(0)p_{0}+a(1)p_{1}}{\max\{a(0),a(1)\}}=\frac{a(1)(p_{0}+p_{1})}{a(1)}=1$
when $a(0)=a(1)$.
However, when $a(0)\neq a(1)$, the estimand’s weights differ from $(p_{0},p_{1})$, the population weights for the two covariate cells. For concreteness, let $a(1)>a(0)$. In this case, the estimand overrepresents the population with $X=1$ and underrepresents the population with $X=0$, relative to the ATE. The largest subpopulation $\{W^{*}=1\}$ that places weights $(a(0),a(1))$ can be constructed by combining subsets of the subpopulations defined by $\{X=0\}$ and $\{X=1\}$. Specifically, let
$W^{*}=\mathbbm{1}(X=1)+\mathbbm{1}\left(U\leq\frac{a(0)}{a(1)},X=0\right),$
where $U\sim\text{Unif}(0,1)$ is independent from $(Y(1),Y(0),X)$. This is a subpopulation that contains all units with $X=1$ and fraction $a(0)/a(1)<1$ of units with $X=0$, selected uniformly at random. Its size is
$\mathbb{P}(W^{*}=1)=\mathbb{P}(X=1)+\mathbb{P}\left(U\leq\frac{a(0)}{a(1)},X=0\right)=p_{1}+\frac{a(0)}{a(1)}p_{0}<1,$
which is the same as $\frac{\mathbb{E}[a(X)]}{\sup_{x\in\{0,1\}}a(x)}=\frac{a(1)p_{1}+a(0)p_{0}}{a(1)}$ obtainable from Theorem 4.1. The average treatment effect in this subpopulation is given by
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{P}(X=0\mid W^{*}=1)\tau_{0}(0)+\mathbb{P}(X=1\mid W^{*}=1)\tau_{0}(1)$
$=\frac{\frac{a(0)}{a(1)}p_{0}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}\tau_{0}(0)+\frac{p_{1}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}\tau_{0}(1)$
$=\frac{a(0)p_{0}}{a(0)p_{0}+a(1)p_{1}}\tau_{0}(0)+\frac{a(1)p_{1}}{a(0)p_{0}+a(1)p_{1}}\tau_{0}(1)$
$=\mu(a,\tau_{0}).$
The relative weights placed on $\{X=0\}$ and $\{X=1\}$ in subpopulation $\{W^{*}=1\}$ are given by
$\frac{\mathbb{P}(X=0\mid W^{*}=1)}{\mathbb{P}(X=1\mid W^{*}=1)}$ $=\frac{\frac{a(0)}{a(1)}p_{0}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}/\frac{p_{1}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}=\frac{\omega_{0}}{\omega_{1}},$
matching the ratio of the weights on $\{X=0\}$ and $\{X=1\}$ assigned by the estimand. The subpopulation $\{W^{*}=1\}$ cannot expand while preserving this ratio since it already encompasses all units with $X=1$. Therefore, $W^{*}$ is the largest subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$.
4.1.2 Partial Identification of Average Effects
This subpopulation size can be used to bound target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ using only the weighted estimand and bounds on $Y$. Given knowledge of the weighted estimand $\mu(a,\tau_{0})$ and its representativeness, given by $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$, we can decompose the target estimand as
$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\cdot\mathbb{P}(W^{*}=1\mid W_{0}=1)$
$\qquad+\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]\cdot(1-\mathbb{P}(W^{*}=1\mid W_{0}=1))$
$=\mu(a,\tau_{0})\cdot\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]\cdot(1-\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$
for $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$. Therefore, $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is point-identified given knowledge of $(\mu(a,\tau_{0}),\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$ if $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=1$, and partially identified if $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})<1$. If we have knowledge of bounds for the treatment effect $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]$, e.g. from the support of the potential outcomes, we can obtain bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. For example, if $\text{supp}(Y(1)-Y(0))=[B_{\ell},B_{u}]$, bounds for the target estimand are given by
$\left[(\mu(a,\tau_{0})-B_{\ell})\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+B_{\ell},(\mu(a,\tau_{0})-B_{u})\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+B_{u}\right].$
The width of these bounds is $(B_{u}-B_{\ell})(1-\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$. Hence for fixed $(B_{\ell},B_{u})$, this width decreases linearly with $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ and values of $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ close to 1, or high degrees of internal validity, lead to narrow bounds.
4.2 Quantifying Internal Validity Given $\tau_{0}$
As in the previous subsection, we can also ask how internally valid a weighted estimand can be, given knowledge of the CATE function. In this case, the object of interest is
$\overline{P}(a,W_{0};\{\tau_{0}\})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1).$ (4.4)
where $\tau_{0}$ is a given CATE function. Since $\tau_{0}$ is known, the condition $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ can be written as
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
or, equivalently,
$\mathbb{E}[(\tau_{0}(X)-\mu(a,\tau_{0}))\underline{w}^{*}(X)\mid W_{0}=1]=0,$ (4.5)
where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$. Equation (4.5) is a linear constraint on the conditional probability of being in subpopulation $W^{*}$. Additionally, the objective function $\mathbb{P}(W^{*}=1\mid W_{0}=1)=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ is linear in $\underline{w}^{*}$. Thus, the optimization in (4.4) can be cast as a linear program. To see this, consider as an example the case where $W_{0}=1$ almost surely and where $X$ is discrete with finite support, i.e. $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$. Let $f_{k}=\mathbb{P}(W^{*}=1,X=x_{k})$ denote the probability of being in subpopulation $W^{*}$ and having covariate value $x_{k}$. Note that $f_{k}\in[0,p_{k}]$ where $p_{k}\equiv\mathbb{P}(X=x_{k})$.
We can write the above optimization problem as
$\max_{(f_{1},\ldots,f_{K})\geq\textbf{0}}\sum_{k=1}^{K}f_{k}\qquad\text{ such that }\qquad$ $f_{k}\leq p_{k}\text{ for }k\in\{1,\ldots,K\},$
$\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))f_{k}=0,$
a finite-dimensional linear program. This program has a feasible solution if $\tau_{0}(x_{k})-\mu(a,\tau_{0})$ is not strictly positive or strictly negative for all $k$, which is precisely stated in the condition for Theorem 3.2. While there exist many methods for solving linear programs, the value function can be obtained through an algorithm that is simple to describe analytically.
Let $\textbf{f}=(f_{1},\ldots,f_{K})$, $\textbf{p}=(p_{1},\ldots,p_{K})$, and $\textbf{t}_{\mu}=(\tau_{0}(x_{1})-\mu(a,\tau_{0}),\ldots,\tau_{0}(x_{K})-\mu(a,\tau_{0}))$. Without loss of generality, assume that $\tau_{0}(x_{1})-\mu(a,\tau_{0})\leq\tau_{0}(x_{2})-\mu(a,\tau_{0})\leq\cdots\leq\tau_{0}(x_{K})-\mu(a,\tau_{0})$.
  1. 1.
[…]
  2. 2.
    If $\textbf{t}_{\mu}^{\prime}\textbf{f}=0$, end the algorithm and report $\sum_{k=1}^{K}f_{k}$.
  3. 3.
    If $\textbf{t}_{\mu}^{\prime}\textbf{f}\neq 0$:
    1. (a)
      If $\textbf{t}_{\mu}^{\prime}\textbf{f}>0$, let $k^{*}=\max\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,\frac{-\sum_{k=1}^{k^{*}-1}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))p_{k}}{\tau_{0}(x_{k^{*}})-\mu(a,\tau_{0})}\right\}$.
    2. (b)
      If $\textbf{t}_{\mu}^{\prime}\textbf{f}<0$, let $k^{*}=\min\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,\frac{-\sum_{k=k^{*}+1}^{K}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))p_{k}}{\tau_{0}(x_{k^{*}})-\mu(a,\tau_{0})}\right\}$.
  4. 4.
    Go to step 2.
When $\mu(a,\tau_{0})$ exceeds $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, this algorithm reduces the weights associated with largest CATEs until $\mu(a,\tau_{0})$ equals $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for some subpopulation. When $\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the same procedure is instead applied to the smallest CATEs. The support assumption of Theorem 3.2 guarantees that this algorithm ends.
When $X$ is not discretely supported, the problem can still be cast as a linear program, but its dimension may be infinite, which generates difficulties in implementation. However, we show this program has an analytical solution for the size of the subpopulation of interest, $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. The following theorem gives its expression and is valid when vector $X$ has an arbitrary kind of distribution, with continuous, discrete, and mixed components, as is often the case in empirical applications.
Theorem 4.2.
Let Assumption 3.1 hold. Let $T_{\mu}=\tau_{0}(X)-\mu(a,\tau_{0})$. If $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})$,
$\overline{P}(a,W_{0};\{\tau_{0}\})$
=$=\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}\\ \qquad\text{ where }\alpha^{+}=\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]\geq 0\}\text{ if }\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}\\ \qquad\text{ where }\alpha^{-}=\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]\leq 0\}\text{ if }\mu(a,\tau_{0})>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]\\ 1\qquad\text{ if }\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1].\end{cases}$ (4.6)
If $\mu(a,\tau_{0})\notin\mathcal{S}(\tau_{0};W_{0})$, then $\overline{P}(a,W_{0};\{\tau_{0}\})=0$.
The computation of these bounds can be done using a linear programming algorithm when $X$ is discrete, or through plug-in estimators of the terms in equation (4.2) regardless of the nature of the support of $X$.
In this setting, the value $\overline{P}(a,W_{0};\{\tau_{0}\})$ is larger when the truncated subpopulations are smaller. In particular, this is the case when there are a few units with extreme values of $\tau_{0}$ whose removal has a large impact on the estimand, but a small impact on the share of the population.
4.2.1 Illustrative Example: A Single Continuous Covariate
To illustrate the previous theorem, let $X$ be a continuously distributed covariate with support $[x_{L},x_{U}]$. Let $W_{0}=1$ almost surely, and let $\mu\equiv\mu(a,\tau_{0})$ denote the estimand. Also, for simplicity assume that $\tau_{0}(x)$ is increasing in $x$ and that $\tau_{0}(x_{L})<\mu<\tau_{0}(x_{U})$. Without loss of generality, let $\mathbb{E}[Y(1)-Y(0)]\geq\mu$. If $\mathbb{E}[Y(1)-Y(0)]=\mu$, then the estimand is perfectly representative of the population since it equals the average treatment effect over it. Now consider the case where $\mathbb{E}[Y(1)-Y(0)]>\mu$. In this case, the estimand is larger than the ATE so it is not representative of the entire population.
We are seeking the largest subpopulation $\{W^{*}=1\}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu$. Equivalently, we can seek the smallest subpopulation $\{W^{-}=1\}$ such that, when it is removed from the original population, the average treatment effect equals $\mu$. The subpopulation $\{W^{-}=1\}$ is related to $\{W^{*}=1\}$ via $W^{-}+W^{*}=1$. We will search for $W^{-}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=0]=\mu$ and $\mathbb{P}(W^{-}=1)$ is minimized.
Since the ATE exceeds $\mu$, we must have that $\mathbb{E}[\tau_{0}(X)\mid W^{-}=1]>\mu$. For a given subpopulation of size $\mathbb{P}(W^{-}=1)$, removing the subpopulation with the largest values of $\tau_{0}(x)$ yields the largest decrease in $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=0]$. Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})$ is obtained by removing a subpopulation of the kind $W^{-}(\delta)=\mathbbm{1}(\tau_{0}(X)\geq\delta)$ for a given threshold $\delta$. This subpopulation consists of units whose CATEs exceed a threshold $\delta$. This threshold is determined by the constraint $\mathbb{E}[\tau_{0}(X)\mid W^{-}(\delta)=0]=\mu$, or
$\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\delta]=\mu.$ (4.7)
This constraint states that once $\{W^{-}(\delta)=1\}$ is removed, the average treatment effect for the remaining units equals the weighted estimand. Since we assumed $\tau_{0}$ is increasing, $W^{*}$ corresponds to all units with covariate values below threshold $\delta=\tau_{0}^{-1}(\delta)$. Therefore,
$W^{*}=\mathbbm{1}(X\leq\tau_{0}^{-1}(\delta)),$
where $\delta$ is the unique solution to (4.7). How large $\mathbb{P}(W^{*}=1)$ is depends on the size of this fraction of removed units. A smaller subpopulation needs to be removed when $\tau_{0}(X)$ has a long right tail, and a larger subpopulation needs to be removed when the distribution of $\tau_{0}(X)$ in this right tail is shorter.
4.3 Generalizing to Subpopulations
The results from the previous two sections can be generalized to cases where we hold the average effect for a subset of $W_{0}$ as the object of interest. Specifically, let $W^{\prime}$ be a regular subpopulation of $W_{0}$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]$ be the target parameter. In analogy to the previous subsections, we ask (i) when does there exist a regular subpopulation $W^{*}$ of $W^{\prime}$ such that $\mu(a,\tau_{0})$ can be written as a causal estimand and (ii) how representative of $W^{\prime}$ is this estimand.
For concreteness, in the IV model of Section 2.2, we may be interested in the average treatment effect for compliers/defiers with covariates in given set $\mathcal{X}_{0}$: $\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X\in\mathcal{X}_{0}]$. In this case $W_{0}=\mathbbm{1}(D(1)>D(0))$ and $W^{\prime}=\mathbbm{1}(D(1)>D(0),X\in\mathcal{X}_{0})$, a regular subpopulation of $W_{0}$.
We first present a result showing conditions for the existence of such population $W^{*}$ in the case where $\tau_{0}$ is unrestricted, and when $\tau_{0}$ is fixed.
Theorem 4.3.
Let Assumption 3.1 hold and let $W^{\prime}$ be a regular subpopulation of $W_{0}$.
  1. 1.
    (Uniformly in $\tau_{0}$) Let $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$. [^5] We use the conventions $0/0=1$ and $a/0=\infty$ when $a>0$. Then, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ if and only if $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$.
  2. 2.
    (Given $\tau_{0}$) There exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$ if and only if $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W^{\prime})$.
Note that letting $W^{\prime}=W_{0}$ yields Theorems 3.1 and 3.2 as special cases since $W_{0}$ is a regular subpopulation of $W_{0}$ by definition.
The condition $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ rules out uniform causal representations if there are covariate values for which $a(x)$ is strictly positive but are not represented in $W^{\prime}$, i.e., $\mathbb{P}(W^{\prime}=1\mid X=x,W_{0}=1)=0$ when $a(x)>0$. For example, if $W^{\prime}$ is a subpopulation of $W_{0}$ defined by a subset of covariate values such that $W^{\prime}=\mathbbm{1}(X\in\mathcal{X}_{0})\cdot W_{0}$, then $W^{\prime}$ is a regular subpopulation but will fail the above inequality if $a(X)>0$ for $X\in\text{supp}(X\mid W_{0}=1)\setminus\mathcal{X}_{0}$. In the case where $\tau_{0}$ is known, the existence of a causal subpopulation of $W^{\prime}$ is equivalent to the estimand being in the convex hull of $\text{supp}(\tau_{0}(X)\mid W^{\prime}=1)$.
Regarding the representativeness of a weighted estimand, we generalize Theorems 4.1 and 4.2 and ask how large $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ can be given that $W^{*}$ is a regular subpopulation of $W^{\prime}$, which is itself a regular subpopulation of $W_{0}$. In what follows, we let $\underline{w}^{\prime}(X)=\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$.
Theorem 4.4.
Let Assumption 3.1 hold and let $W^{\prime}$ be a regular subpopulation of $W_{0}$.
  1. 1.
    (Uniformly in $\tau_{0}$) Let $\sup\left(\text{supp}\left(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1\right)\right)<\infty$. If $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$, then
    $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})=\mathbb{E}[a(X)\mid W_{0}=1]\cdot\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}\cdot\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right).$ (4.8)
    If $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$, $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})=0$.
  2. 2.
    (Given $\tau_{0}$) If $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W^{\prime})$, then
    $\overline{P}(a,W^{\prime};\{\tau_{0}\})$
    =$=\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W^{\prime}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W^{\prime}=1]}{\alpha^{+}}\\ \qquad\text{ where }\alpha^{+}=\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W^{\prime}=1]\geq 0\}\text{ if }\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W^{\prime}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W^{\prime}=1]}{\alpha^{-}}\\ \qquad\text{ where }\alpha^{-}=\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]\leq 0\}\text{ if }\mu(a,\tau_{0})>\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]\\ 1\qquad\text{if }\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1].\end{cases}$ (4.9)
    If $\mu(a,\tau_{0})\notin\mathcal{S}(\tau_{0};W^{\prime})$, then $\overline{P}(a,W^{\prime};\{\tau_{0}\})=0$.
We can also use this result to obtain bounds on $\mathbb{P}(W^{*}=1)$, or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. This can be done by noting that $W^{*}$ is a subpopulation of both $W^{\prime}$ and $W_{0}$, so
$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\cdot\mathbb{P}(W^{\prime}=1)\quad\text{ and }\quad\mathbb{P}(W^{*}=1\mid W_{0}=1)=\frac{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\cdot\mathbb{P}(W^{\prime}=1)}{\mathbb{P}(W_{0}=1)}.$
Thus bounds on these probabilities are trivially obtained from the bound on $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ from Theorem 4.4 and knowledge of $(\mathbb{P}(W^{\prime}=1),\mathbb{P}(W_{0}=1))$.
4.4 Examples Revisited
For simplicity, we assume that $\sup(\text{supp}(a(X)\mid W_{0}=1))=\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)$ for the remainder of the paper. This condition is satisfied for example when $a(\cdot)$ is continuous.
4.4.1 Unconfoundedness
Recall the expression for the coefficient on $D$ in a population regression of $Y$ on $(1,D,X)$:
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[p(X)(1-p(X))\tau_{0}(X)]}{\mathbb{E}[p(X)(1-p(X))]}.$
Suppose the target estimand is the average treatment effect, i.e. $W_{0}=1$ almost surely.
By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function $a(X)\equiv p(X)(1-p(X))$ is non-negative. By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is given by
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[p(X)(1-p(X))]}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}.$
The corresponding subpopulation $W^{*}$ satisfies
$\mathbb{P}(W^{*}=1\mid X)$ $=\frac{p(X)(1-p(X))}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}.$
This is a subpopulation that places greater weights on units which have a larger variation in treatment given their covariate values. The size of this subpopulation is largest when $\operatorname{var}(D\mid X)=p(X)(1-p(X))$ is constant, in which case $\mathbb{P}(W^{*}=1\mid X)=1$. This is the case if and only if $p(X)$ has support equal to $\{b,1-b\}$ for some $b\in[0,1]$. This is implied by $D\perp\!\!\!\perp X$, or random assignment. It can also be achieved if there exists a partition of $\text{supp}(X)$ where $\mathbb{P}(D=1\mid X)=b$ on one element and $\mathbb{P}(D=1\mid X)=1-b$ on its complement. Whenever $\operatorname{var}(p(X)(1-p(X)))>0$, $\{W^{*}=1\}$ will be a strict subpopulation.
The size of this subpopulation is the expectation of $\operatorname{var}(D\mid X)$ divided by its maximum value. There are a few ways this expression can be further simplified or bounded. Its numerator is bounded above by $\operatorname{var}(D)=\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)$, which is particularly simple to estimate. As for the denominator, it is a non-smooth functional of $p(\cdot)$ which will lead to nonstandard inference. However, if $X$ is continuously distributed, it may be likely that $p(X)$ is continuously distributed as well. In this case, it is likely that $1/2\in\text{supp}(p(X))$. If this is the case, $\sup_{x\in\text{supp}(X)}p(x)(1-p(x))=1/4$. Combining these approximations yields that
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $\leq 4\cdot\mathbb{P}(D=1)\cdot\mathbb{P}(D=0),$
when the support of $p(X)$ includes 1/2. This bound is trivial when $\mathbb{P}(D=1)=1/2$, but is informative when the treatment probability is close to 0 or 1. For example, if $\mathbb{P}(D=1)=0.1$, the OLS estimand cannot causally represent more than 36% of the population.
When $1/2\in\text{supp}(p(X))$, we can also compute bounds on the ATE derived from the OLS estimand, bounds on the support of $(Y(1),Y(0))$, and our measure of representativeness $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$. Following Remark 4.1.2, bounds on the ATE are given by
$\left[(\beta_{\text{OLS}}-B_{\ell})4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{\ell},(\beta_{\text{OLS}}-B_{u})4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{u}\right].$
Estimating these bounds requires the estimation of one additional quantity beyond the OLS estimator, which is the expectation of $\operatorname{var}(D\mid X)$. Of course, these bounds’ width depends crucially on $B_{u}-B_{\ell}$, or the width of the support for unit-level treatment effects.
Alternatively, we can assess the representativeness of $\beta_{\text{OLS}}$ with respect to an alternative estimand such as $\mathbb{E}[Y(1)-Y(0)\mid D=1]$, the average treatment effect on the treated. In this case, we can write
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]},$
where $w_{0}(X)=\mathbb{P}(D=1\mid X)=p(X)$, the propensity score, and $a(X)=\mathbb{P}(D=0\mid X)=1-p(X)$. Applying Theorem 4.1 (or 4.4) yields that
$\overline{P}(a,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[1-p(X)\mid D=1]}{\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)\cdot\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}$
is the largest value that $\mathbb{P}(W^{*}=1\mid D=1)$ can take. Once again, this bound depends only on the propensity score and the distribution of $X$. The same bound is obtained by using Theorem 4.4 instead and setting $W_{0}=1$ and $W^{\prime}=D$. This subpopulation satisfies
$\mathbb{P}(W^{*}=1\mid X,D=1)=\frac{1-p(X)}{1-\inf_{x\in\text{supp}(X\mid D=1)}p(X)}$
so units with smaller propensity scores are more likely to be included in $W^{*}$, given that they are treated. This probability is maximized at 1 when $p(X)$ is constant, or if $D\perp\!\!\!\perp X$. In this case, $\mathbb{P}(W^{*}=1\mid D=1)=1$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(D=1)$.
If $p(X)$ takes values close to 0, this bound equals
$\overline{P}(a,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)}\leq\frac{\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)}{\mathbb{P}(D=1)}=\mathbb{P}(D=0).$
This suggests that the OLS estimand is more representative of the ATT when the fraction of untreated units is larger.
We can also assess the representativeness of $\beta_{\text{OLS}}$ given $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. For simplicity, assume that $X$ has a continuous distribution and, without loss of generality, assume that $\beta_{\text{OLS}}>\text{ATE}$. Then, using Theorem 4.2, we obtain
$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\mathbb{P}(\tau_{0}(X)\leq b^{*}),$
where $b^{*}$ satisfies $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq b^{*}]=\beta_{\text{OLS}}$. The quantity $\overline{P}(a,W_{0};\{\tau_{0}\})$ is largest when the least amount of trimming needs to be applied. This is the case when the trimmed values are largest, or when $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\geq b]$ is large when $b$ is near the maximum of $\text{supp}(\tau_{0}(X))$. More concretely, if some covariate values in $\text{supp}(X)$ have large CATEs, then considering a subpopulation that removes only a small subset of $\text{supp}(X)$ — those corresponding to large CATE values — can allow $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq b^{*}]$ and $\beta_{\text{OLS}}$ to be equal.
4.4.2 Instrumental Variables
Consider again the setting of Section 2.2. The estimand $\beta_{\text{2SLS}}$ can be characterized as $\mu(a_{\text{2SLS}},\tau_{0})$ for $a_{\text{2SLS}}(X)=\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$.
Since $a_{\text{2SLS}}(X)\geq 0$, there exists a subpopulation of $\{W_{0}=1\}$ such that $\beta_{\text{2SLS}}$ is an average treatment effect over that subpopulation. The maximum size of that subpopulation is given by
$\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{2SLS}}(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}$
$=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}{\mathbb{P}(D(1)\neq D(0))\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\operatorname{var}(Z\mid X=x)\cdot\mathbb{P}(D(1)\neq D(0)\mid X=x)}.$
Under Assumptions 2.2 and 2.4, $\mathbb{P}(D(1)\neq D(0)\mid X)$ is identified from the distribution of $(D,X,Z)$ as $w_{0}(X)\equiv|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|$, which is the fraction of compliers or defiers, i.e. units for which $D(1)\neq D(0)$, among those with covariates $X$.
The maximum value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is obtained when $\operatorname{var}(Z\mid X)\cdot|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|=|\operatorname{cov}(Z,D\mid X)|$ does not depend on $X$. This occurs, for example, when the instrument and the fraction of units for which $D(1)\neq D(0)$ are independent of $X$. In this case, we have that $\beta_{\text{2SLS}}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average effect of treatment in the complier/defier subpopulation.
The IV estimand has the same $W_{0}$, $w_{0}(X)$, but has $a_{\text{IV}}(X)=\text{sign}(\mathbb{P}(D(1)\geq D(0)\mid X)-\mathbb{P}(D(1)\leq D(0)\mid X))\cdot\operatorname{var}(Z\mid X)\equiv c(X)\operatorname{var}(Z\mid X)$ instead.
First, we notice that if $\mathbb{P}(c(X)=-1)>0$, then weights are negative with positive probability and there does not exist a causal representation for the estimand $\beta_{\text{IV}}$ that is uniform in $\tau_{0}\in\mathcal{T}_{\text{all}}$. However, there will exist a causal representation given $\tau_{0}$ if the support condition of Theorem 3.2 holds, in that $\beta_{\text{IV}}$ lies in the support of $\tau_{0}(X)$.
If we assume strong monotonicity (Assumption 2.3), then $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)\geq 0$ and
$\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{IV}}(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}$
$=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)>D(0)\mid X)]}{\mathbb{P}(D(1)>D(0))\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\operatorname{var}(Z\mid X=x)}.$
Here the representativeness of the IV estimand is maximized when $\operatorname{var}(Z\mid X)$ is constant, which occurs when $Z$ is independent of $X$. In this case, $\beta_{\text{IV}}$ equals LATE. The quantities $\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ and $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ are not ranked uniformly in the distributions of $(D(1),D(0),X,Z)$ as there are data-generating processes that make each of these two quantities larger than the other. For example, if $\operatorname{var}(Z\mid X)$ is constant but $\mathbb{P}(D(1)\neq D(0)\mid X)$ is not, then $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})<\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$. This scenario is plausible if $Z$ is randomly assigned and $X$ is a vector of pre-assignment characteristics. This inequality is reversed if $a_{\text{2SLS}}(X)=|\operatorname{cov}(Z,D\mid X)|$ is constant but $\mathbb{P}(D(1)\neq D(0)\mid X)$ is not. They are equally representative when $\mathbb{P}(D(1)\neq D(0)\mid X)$ is constant. In this case, the estimands are equal, so this is not unexpected.
4.4.3 Differences-in-Differences
We now consider the weights obtained in Proposition 2.5 under its assumptions. These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. Using Theorem 4.1, the level of internal validity of $\beta_{\text{TWFE}}$ relative to $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
$\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{B}(G)w_{0}(G)]}{\mathbb{E}[w_{0}(G)]\cdot\sup_{g\in\text{supp}(G\mid D=1)}a_{B}(g)}$
$=\frac{\sum_{g=2}^{T}\operatorname{var}(D\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\cdot\mathbb{P}(G=g)}{\mathbb{P}(D=1)\cdot\sup_{g\in\{2,\ldots,T\}}\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))}.$
Due to the absorbing nature of treatment in our setting, all expressions involving the distribution of $D$ given $P$ or $G$ can be derived as a function of the marginal distribution of $G$. Therefore, $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ depends only on $\{\mathbb{P}(G=g)\}_{g\in\{2,\ldots,T\}}$.
To give some intuition, consider the case where $T=3$ and therefore $G\in\{2,3,+\infty\}$. In this case, $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ will simply be a function of $(\mathbb{P}(G=2),\mathbb{P}(G=3))$.
Calculations yield that
$\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ =$=\begin{cases}\frac{2\mathbb{P}(G=2)+\omega\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}&\text{ if }\omega<1\\ \frac{\omega^{-1}2\mathbb{P}(G=2)+\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}&\text{ if }\omega>1\\ 1&\text{ if }\omega=1,\end{cases}$
where
$\omega$ $=\frac{4-2\mathbb{P}(G=2)-4\mathbb{P}(G=3)}{2-2\mathbb{P}(G=2)-\mathbb{P}(G=3)}=\frac{a_{B}(2)}{a_{B}(3)}.$
Therefore, the TWFE estimand is perfectly representative of the ATT if and only if $\omega=1$.
With some algebra, we can see that $\omega<1$ if and only if $\mathbb{P}(G=3)>2/3$, while $\omega>1$ if and only if $\mathbb{P}(G=3)<2/3$. We can also see that $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ is equal to 1 when $\mathbb{P}(G=3)=2/3$, and that the representativeness of the TWFE estimand declines as $|\mathbb{P}(G=3)-2/3|$ increases. This is due to weight function $a_{B}(g)$ being constant in $g$ if and only if $\mathbb{P}(G=3)$ equals 2/3. Whenever the weight function is constant, we have that $\mu(a,\tau_{0})=\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, and therefore the weighted estimand equals the average effect over population $\{W_{0}=1\}$.
5 Discussion of Estimation and Inference
We now briefly discuss procedures to estimate $\overline{P}(a,W_{0};\mathcal{T})$.
First consider the case when $\mathcal{T}=\mathcal{T}_{\text{all}}$. Then, we seek to estimate
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}=\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}.$ (5.1)
Suppose we observe a random sample of size $n$, $\{(W_{i},X_{i})\}_{i=1}^{n}$, where $W_{i}$ are a set of variables that allow the estimation of $a(\cdot)$ and $w_{0}(\cdot)$. For example, under unconfoundedness we can let $W_{i}=(Y_{i},D_{i})$ or $W_{i}=D_{i}$ since only the distribution of $(D,X)$ is needed to identify $a(\cdot)$. In our instrumental variables examples, one can let $W_{i}=(Z_{i},D_{i})$.
Assuming the existence of estimators for $a(\cdot)$ and $w_{0}(\cdot)$, we can propose the following analog estimator:
$\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)}.$
This estimator requires knowledge of the support of $X$ given $W_{0}=1$, but can be implemented by taking the maximum over the empirical distribution of $X\mid W_{0}=1$. Consistency of $\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$ for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is achieved when
[…]
and $\widehat{a}(x)$ is uniformly consistent for $a(x)$ over $x\in\text{supp}(X\mid W_{0}=1)$. These conditions can be achieved under many standard estimators of $a(\cdot)$ and $w_{0}(\cdot)$. For example, when $a(X)$ and $w_{0}(X)$ consist of continuous functions of conditional means, as is the case when $a(X)=\operatorname{var}(D\mid X)$ or, for the IV estimator, $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ and $w_{0}(X)=|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|$, consistent estimators of these conditional means will result in the consistency of $\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$. Cell-by-cell estimators can be used when $X_{i}$ is discrete, or many nonparametric estimators when $X_{i}$ contains continuous components. See Hansen (2022) for a recent textbook treatment of such methods.
Inference will generally be nonstandard. This is because the asymptotic behavior of the denominator will usually be non-Gaussian due to the lack of Hadamard differentiability of the supremum operator. This lack of differentiability will invalidate standard bootstrap inference: see Fang and Santos (2019). However, the supremum is a Hadamard directionally differentiable functional and nonstandard bootstrap approaches, such as those proposed in Hong and Li (2018) and Fang and Santos (2019), can be used.
We can also view these inferential problems through the lens of intersection or union bounds. For example, we can write
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\inf_{x\in\text{supp}(X\mid W_{0}=1)}\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a(x)}\equiv\inf_{x\in\text{supp}(X\mid W_{0}=1)}\overline{P}(x).$
Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $[0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds. Chernozhukov, Lee, and Rosen (2013) offer methods for such problems. Equivalently, the computation of a one-sided confidence interval $[\widehat{\overline{P}}^{-},1]$ is related to inferential questions in union bounds: see Bei (2024). We leave all details for future work.
Estimation and inference of $\overline{P}(a,W_{0};\{\tau_{0}\})$ is related to the question of estimation and inference in linear programs with estimated constraints. See Fang, Santos, Shaikh, and Torgovitsky (2023) and Cho and Russell (2024) for recent advances on this topic.
6 Empirical Application
In this section, we implement the proposed tools in an application to the effects of unilateral divorce laws in the U.S. on female suicide, as in Stevenson and Wolfers (2006). Between 1969 and 1985, 37 states (including the District of Columbia) reformed their law by enabling each spouse to seek divorce without the other spouse’s consent. Stevenson and Wolfers (2006) argue that these “unilateral” or “no-fault” divorce laws reduced female suicide, domestic violence, and spousal homicide. The results on female suicide are also replicated by Goodman-Bacon (2021), whose analysis we follow here.
Our sample consists of 41 states observed over the 1964–1996 period. The outcome of interest is the state- and year-specific female suicide rate (per million women), as computed by the National Center for Health Statistics (NCHS). The treatment is whether the state allowed unilateral divorce in a given year. Following Goodman-Bacon (2021), our sample omits Alaska and Hawaii. Additionally, we omit Louisiana, Maryland, North Carolina, Oklahoma, Utah, Vermont, Virginia, and West Virginia. These eight states (and Alaska) had unilateral divorce laws preceding 1964 and are therefore always treated within our timeframe.
{adjustwidth}
-1in-1in A. Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B. Representativeness of the TWFE estimand based on Proposition 2.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C. Representativeness of the TWFE estimand based on Proposition 2.5 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
Table 1: Representativeness of the TWFE estimand of the effects of unilateral divorce laws
Panel A of Table 1 reports our baseline estimates of the average effects of unilateral divorce laws on female suicide. After we drop the eight always-treated states, the TWFE estimate, –0.604, becomes much smaller in absolute value than the corresponding estimate in Goodman-Bacon (2021), –3.080. Unlike that estimate, ours is also statistically insignificant, with $p$-value = 0.819.
The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified. Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138. The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001. As is clear, the choice of an estimation approach has fundamental implications for our conclusions in this study.
While the TWFE estimate and both estimates of the ATT are clearly different, this paper focuses on another implication of the nonuniformity of the TWFE weight function. We ask: How representative of the underlying population is the TWFE estimand? What is the internal validity of this estimand if we are interested in the treated subpopulation? Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), restated in our Proposition 2.4. First, because the weights on some of the group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$. Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
Panel C of Table 1 revisits these questions on the basis of the representation of the TWFE estimand in Proposition 2.5, based on Goodman-Bacon (2021). Here, we assume that group-level average treatment effects are constant over time, which eliminates the problem of negative weights. Indeed, we now conclude that the TWFE estimand has a causal interpretation uniformly in $\tau_{0}$, even if it is still not particularly representative of the underlying population or the treated subpopulation. Our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$ are equal to 14.00% and 22.46%, respectively. When we use the estimated CATE function in constructing the bounds, these estimates increase to 48.02% and 77.07%. This is obviously much more than our initial estimate of 0, but still substantially less than 1, guaranteed in the case of $\mathbb{P}(W^{*}=1\mid D=1)$ when using the estimation approaches in Callaway and Sant’Anna (2021), Wooldridge (2021), and other recent papers.
7 Conclusion
In this paper, we studied the representativeness and internal validity of a class of weighted estimands, which includes the popular OLS, 2SLS, and TWFE estimands in additive linear models. We examined the conditions under which such estimands can be written as the average treatment effect for some (possibly latent) subpopulation. In our main result, we derived the sharp upper bound on the size of that subpopulation. We consider this bound to be a valuable diagnostic for empirical research. When a given estimand can be shown to correspond to the average treatment effect for a large subset of the population of interest, we say its internal validity is high. In an application to the effects of unilateral divorce laws in the U.S. on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021), we showed that the TWFE estimand has a low degree of internal validity (assuming that the treated subpopulation is of interest), even when we assume away the existence of negative weights. Because this result is then necessarily driven by the nonuniformity of the TWFE weight function, it corroborates the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
Appendix
This appendix is organized as follows: Appendix A contains proofs for Section 2, Appendix B contains proofs for Section 3, Appendix C contains proofs for Section 4, and Appendix D contains additional results and derivations regarding difference-in-differences and associated weighted estimands.
Appendix A Proofs for Section 2
Proof of Proposition 2.4.
We begin by noting that
$\beta_{\text{TWFE}}$ $=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}]}$
$=\frac{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}\mid P=t]\mathbb{P}(P=t)}{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}\mid P=t]\mathbb{P}(P=t)}$
$=\frac{\mathbb{E}[\ddot{D}Y]}{\mathbb{E}[\ddot{D}^{2}]},$
where the second equality follows from the uniform distribution of $P$ which is independent from $(\ddot{D}_{t},Y_{t})$ for all $t\in\{1,\ldots,T\}$. The third equality follows from defining $\ddot{D}\equiv\ddot{D}_{P}$. We also note that
$\ddot{D}_{P}$ $=D_{P}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\sum_{t=1}^{T}\mathbb{E}[D_{t}]\mathbbm{1}(P=t)+\sum_{s=1}^{T}\mathbb{E}[D_{s}]\mathbb{E}[\mathbbm{1}(P=s)]$
$=D-\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)-\mathbb{E}[D\mid P]+\mathbb{E}[D]$
$=D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D].$
The third equality follows from $\mathbb{E}[D\mid G]=\mathbb{E}[\mathbbm{1}(G\leq P)\mid G]=\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)=\frac{1}{T}\sum_{s=1}^{T}D_{s}$.
We break down the rest of this proof into four steps.
Step 1: Splitting the Numerator in Two
We have that
$\mathbb{E}[\ddot{D}Y]=\mathbb{E}[\ddot{D}(Y(0)+D(Y(1)-Y(0)))]=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid P,G]]$
The first equality follows from $Y=Y(0)+D(Y(1)-Y(0))$ and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
Step 2: First Numerator Term
We have that
$\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]$ $=\mathbb{E}[\ddot{D}\theta(G,P)]=\mathbb{E}[\ddot{D}\ddot{\theta}(G,P)]$
where $\theta(G,P)=\mathbb{E}[Y(0)\mid G,P]$. The second equality follows by properties of projections and from defining $\ddot{\theta}(G,P)$ as follows:
$\ddot{\theta}(G,P)$ $\equiv\theta(G,P)-\mathbb{E}[\theta(G,P)\mid G]-\mathbb{E}[\theta(G,P)\mid P]+\mathbb{E}[\theta(G,P)]$
$=\mathbb{E}[Y(0)\mid G,P]-\mathbb{E}[Y(0)\mid G]-\mathbb{E}[Y(0)\mid P]+\mathbb{E}[Y(0)].$
Then, we note that
$\theta(g^{\prime},t^{\prime})$ $=\mathbb{E}[Y(0)\mid G=g^{\prime},P=t^{\prime}]-\mathbb{E}[Y(0)\mid G=g^{\prime}]-\mathbb{E}[Y(0)\mid P=t^{\prime}]+\mathbb{E}[Y(0)]$
$=\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g^{\prime}]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g).$
Assumption 2.6.2 implies that for any pair $t,t^{\prime}\in\{1,\ldots,T\}$ and any $g^{\prime}\in\mathcal{G}$
$\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]=\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)].$
This can be shown for $t^{\prime}>t$ by writing $\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]=\sum_{s=t+1}^{t^{\prime}}\mathbb{E}[Y_{s}(0)-Y_{s-1}(0)\mid G=g^{\prime}]=\sum_{s=t+1}^{t^{\prime}}\mathbb{E}[Y_{s}(0)-Y_{s-1}(0)]=\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]$. Similar derivations show this holds for $t^{\prime}<t$. The case where $t^{\prime}=t$ is trivial. Therefore,
$\theta(g^{\prime},t^{\prime})$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]\right)\mathbb{P}(G=g)$
$=0$
for all $(g^{\prime},t^{\prime})\in\mathcal{G}\times\{1,\ldots,T\}$, which implies $\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]=0$.
Step 3: Second Numerator Term
We can write
$\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[D(Y(1)-Y(0))\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)].$
The first equality is by definition, the second by $D^{2}=D$, the third by $\mathbb{E}[D\mid G,P]=D$, and the fourth by law of total probability.
Step 4: Denominator
In this step, we show that
$\mathbb{E}[\ddot{D}^{2}]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)].$
The first line is obtained from properties of linear projections, the second from $D^{2}=D$, and the third from $D=\mathbb{P}(D=1\mid G,P)$.
We conclude the proof by noting the equivalence of integrating over the distribution of $P$ and averages over time periods, which shows the equivalence between
$\beta_{\text{TWFE}}=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)]}$
and the expression in Proposition 2.4. ∎
Proof of Proposition 2.5.
Proposition 2.4 and $\mathbb{P}(D=1\mid G,P)=D$ yielded
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]}.$
Since $\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]=\mathbb{E}[Y(1)-Y(0)\mid G,D=1]$ by assumption, we can use the law of iterated expectations to obtain
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]]}.$
We now calculate the conditional expectation $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$ for $g\in\mathcal{G}$. If $g=+\infty$, then this conditional expectation is 0 by construction, so we let $g\in\{2,\ldots,T\}$ in what follows:
$\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[D\mid G=g]-\frac{\mathbb{E}[\mathbb{E}[D\mid P]D\mid G=g]}{\mathbb{E}[D\mid G=g]}+\mathbb{E}[D]\right)$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(G\leq P)\mid G=g]}{\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]}+\mathbb{E}[\mathbb{E}[\mathbbm{1}(G\leq P)\mid P]]\right)$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(g\leq P)]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(g\leq P)]}{\mathbb{E}[\mathbbm{1}(g\leq P)]}+\mathbb{E}[F_{G}(P)]\right)$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(g\leq P)]-\mathbb{E}[F_{G}(P)\mid g\leq P]\right.$
$\quad\left.+\;\mathbb{E}[F_{G}(P)\mid g\leq P]\mathbb{E}[\mathbbm{1}(g\leq P)]+\mathbb{E}[F_{G}(P)\mid g>P]\mathbb{E}[\mathbbm{1}(g>P)]\right)$
$=\mathbb{E}[D\mid G=g]\mathbb{E}[\mathbbm{1}(g>P)](1+\mathbb{E}[F_{G}(P)\mid g>P]-\mathbb{E}[F_{G}(P)\mid g\leq P])$
$=\mathbb{E}[D\mid G=g](1-\mathbb{E}[D\mid G=g])(1+\mathbb{E}[D\mid g>P]-\mathbb{E}[D\mid g\leq P])$
$=\mathbb{P}(D=1\mid G=g)\mathbb{P}(D=0\mid G=g)(\mathbb{P}(D=1\mid P<g)+\mathbb{P}(D=0\mid P\geq g)).$
The first equality follows from $\mathbb{E}[D\mid G=g]>0$ for $g\in\{2,\ldots,T\}$. The second follows from $D=\mathbbm{1}(G\leq P)$ and the law of iterated expectations, the third from $G\perp\!\!\!\perp P$, the fourth from definitions of conditional expectations and the law of iterated expectations, the fifth from combining terms, the seventh from the law of iterated expectations again, and the last line is obtained by the fact that $D\in\{0,1\}$. The representation in Proposition 2.5 follows. ∎
Appendix B Proofs for Section 3
Proof of Proposition 3.1.
We begin by showing the first claim of the proposition. The equation
$\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ (B.1)
holds since $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, which holds by Definition 3.1. Since $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is assumed positive, we can divide both sides of equation (B.1) by it and obtain
$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ $=\frac{\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]}{\mathbb{P}(W^{*}=1\mid W_{0}=1,X)}$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1,X]$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X],$
where the second equality holds by definition, and the third holds from $W^{*}$ being a subpopulation of $W_{0}$.
We now show the second claim of the proposition. Equation (3.4) can be obtained as follows:
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1]$
$=\frac{\mathbb{E}[W^{*}(Y(1)-Y(0))\mid W_{0}=1]}{\mathbb{E}[W^{*}\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\mathbb{E}[W^{*}(Y(1)-Y(0))\mid X,W_{0}=1]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\mu(\underline{w}^{*},\tau_{0}).$
The first equality follows from $W^{*}$ being a subpopulation of $W_{0}$, the second from the definition of conditional expectation and $\mathbb{P}(W^{*}=1)>0$, the third from the law of iterated expectations and $W^{*}=1$ implying $W_{0}=1$, the fourth from $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, and the fifth and sixth follow immediately. ∎
Proof of Lemma 3.1.
We verify that $W^{*}$ satisfies the two conditions in Definition 3.1. Condition 1 holds since $\mathbb{P}(W^{*}=1\mid W_{0}=0)=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid W_{0}=1)\leq\mathbb{P}(W^{\prime}=1\mid W_{0}=1)=0$, The first equality follows from $W^{*}\leq W^{\prime}$ almost surely, and the last one from $W^{\prime}\in\text{SP}(W_{0})$.
To verify condition 2, note that
$\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W_{0}=1)$ $=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W_{0}=1).$
The first and seventh line follow from $W^{*}\leq W^{\prime}$ almost surely. The second and sixth line follow from factoring conditional probabilities. The third and fifth line follow from $W^{\prime}\leq W_{0}$ almost surely. The fourth line follows from $W^{\prime}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$ and $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$. ∎
Proof of Theorem 3.1.
This theorem follows as a special case of the first part of Theorem 4.3 when $W^{\prime}=W_{0}$. This is because $W_{0}$ trivially regular subpopulation of $W_{0}$, and because the condition
$\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$
is equivalent to $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. They are equivalent since $\underline{w}^{\prime}(X)=\mathbb{P}(W_{0}=1\mid X,W_{0}=1)=1$ almost surely. ∎
Proof of Theorem 3.2.
This is a special case of the second part of Theorem 4.3 where we set $W^{\prime}=W_{0}$. ∎
Proof of Proposition 3.2.
First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. Therefore, by Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$. Therefore,
00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[a(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=d^{\prime}\left(\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\right)$
for all $d\in\mathbb{R}^{d_{X}}$, which implies that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[w^{*}(X)\mid W_{0}=1]}$. Letting $u(x)=\underline{w}^{*}(x)/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we have that
[…]
$W^{*}=\mathbbm{1}\left(U\leq\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\right)\cdot W_{0}$
where $U\perp\!\!\!\perp(X,Y(1),Y(0),W_{0})$. Then $W^{*}$ is a regular subpopulation of $W_{0}$ and $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)=\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}$ since $\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\in[0,1]$ with probability 1 given $W_{0}=1$. Therefore, for all $\tau_{0}(x)=c+d^{\prime}x\in\mathcal{T}_{\text{lin}}$ we have that
$\mu(\underline{w}^{*},\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[u(X)\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[u(X)(c+d^{\prime}X)\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
$=c+d^{\prime}\frac{\mathbb{E}\left[u(X)X\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
$=c+d^{\prime}\frac{\mathbb{E}\left[a(X)X\mid W_{0}=1\right]}{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[a(X)(c+d^{\prime}X)\mid W_{0}=1\right]}{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}$
$=\mu(a,\tau_{0}).$
Therefore, by Proposition 3.1 we have that $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. ∎
Proof of Proposition 3.3.
We consider the $K>0$ case first and the $K=0$ case second.
Case 1: $K>0$.
First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))=\emptyset$ by way of contradiction.
Suppose there is a $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ and let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\in[0,1]$. Let $\tau^{*}(x)=K\cdot\mathbbm{1}(a(x)<0)$. This definition implies $\tau^{*}\in\mathcal{T}_{\text{BV}}(K)$. Since $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ we must have that $\mu(a,\tau_{0})=\mu(\underline{w}^{*},\tau_{0})$ for all $\tau_{0}\in\mathcal{T}_{\text{BV}}(K)$. Since $\tau^{*}\in\mathcal{T}_{\text{BV}}(K)$,
00 0 $=\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})$
$=\mathbb{E}\left[\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\tau^{*}(X)\mid W_{0}=1\right]$
$=K\cdot\mathbb{E}\left[\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\mathbbm{1}(a(X)<0)\mid W_{0}=1\right].$ (B.2)
The right-hand side of (B.2) is $K$ times the expectation of a nonpositive function. This follows from $a(X)\mathbbm{1}(a(X)<0)\leq 0$, $\underline{w}^{*}(X)\mathbbm{1}(a(X)<0)\geq 0$ by $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\geq 0$, $\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1, and $\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]=\mathbb{P}(W^{*}=1\mid W_{0}=1)>0$ by $W^{*}\in\text{SP}(W_{0})$. Since $K>0$, equation (B.2) implies the nonpositive function must equal 0 with probability 1 given $W_{0}=1$:
11 11 $=\mathbb{P}\left(\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\mathbbm{1}(a(X)<0)=0\mid W_{0}=1\right)$
$=\mathbb{P}\left(a(X)=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)]}\cdot\underline{w}^{*}(X)=0\mid a(X)<0,W_{0}=1\right),$
where the second equality follows from $\mathbb{P}(a(X)<0\mid W_{0}=1)>0$. This implies $a(X)$ equals a positive multiple of $\underline{w}^{*}(X)$, a nonnegative quantity, with probability 1 given $\{a(X)<0,W_{0}=1\}$, an event with positive probability that implies $a(X)$ is strictly negative. This is a contradiction and therefore $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))=\emptyset$.
Second, suppose $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$. By Theorem 3.1, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})\neq\emptyset$. Since $\mathcal{T}_{\text{BV}}(K)\subseteq\mathcal{T}_{\text{all}}$, we have that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})\subseteq\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$. Therefore, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))\neq\emptyset$ thus $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{BV}}(K)$.
Case 2: $K=0$.
When $K=0$, the function class $\mathcal{T}_{\text{BV}}(K)$ is the set of all constant functions. In this case, $\tau_{0}(X)=t_{0}$, where $t_{0}\in\mathbb{R}$ denotes a constant. Thus $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(0))\neq\emptyset$ for all weight functions $a(\cdot)$ since $W_{0}\in\text{SP}(W_{0})$ and because $\mu(a,\tau_{0})=\mathbb{E}[a(X)t_{0}\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=t_{0}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ for all $a(\cdot)$. ∎
Appendix C Proofs for Section 4
Proof of Theorem 4.1.
This is a special case of the first part of Theorem 4.4 where we set $W^{\prime}=W_{0}$. ∎
We begin with a technical lemma that we use to prove the subsequent theorems.
Lemma C.1.
Let Assumption 3.1 hold. Let $T_{\mu}=\tau_{0}(X)-\mu$. Then, for any $W^{\prime}\in\text{SP}(W_{0})$, we have that
  1. 1.
    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]$ are left-continuous.
  2. 2.
    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}>\alpha)\mid W^{\prime}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W^{\prime}=1]$ are right-continuous.
Proof of Lemma C.1.
The function $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ is left-continuous if for any strictly increasing sequence $\alpha_{n}\uparrow\alpha$ we have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W^{\prime}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$. To see this holds, note that $f_{n}(t)\equiv t\mathbbm{1}(t<\alpha_{n})\rightarrow t\mathbbm{1}(t<\alpha)$ since $t\mathbbm{1}(t<\alpha_{n})=0$ for all $t\geq\alpha$, and $t\mathbbm{1}(t<\alpha_{n})=t$ whenever $t<\alpha$ for sufficiently large $n$. The random variable $|T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})|$ is dominated by $|T_{\mu}|$ and $\mathbb{E}[|T_{\mu}|\mid W^{\prime}=1]<\infty$ by Assumption 3.1 and by $\mathbb{P}(W^{\prime}=1)>0$. Therefore, by dominated convergence, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W^{\prime}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ hence $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ is left-continuous. The function $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]$ is also left-continuous because $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]=\mathbb{E}[T_{\mu}\mid W^{\prime}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$.
The lemma’s second claim can be similarly shown by considering a sequence $\alpha_{n}\downarrow\alpha$. ∎
Proof of Theorem 4.2.
To simplify the notation in the proof, let $\mu\equiv\mu(a,\tau_{0})$. We break down this proof into four cases.
Case 1: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
We want to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ over the subpopulations $W^{*}$ in $\mathcal{W}(a;W_{0},\{\tau_{0}\})$. Recall that $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if $\mu=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$ and $W^{*}\in\text{SP}(W_{0})$ hold, where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$.
Therefore,
$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\max_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
$=\max_{W^{*}\in\{0,1\}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],W^{*}\in\text{SP}(W_{0})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
$\leq\max_{W^{*}\in\{0,1\}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]$
$=\max_{\underline{w}^{*}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
We will first show a closed-form expression for an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. Then, we will show that this upper bound can be attained by a corresponding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, and therefore it equals $\overline{P}(a,W_{0};\{\tau_{0}\})$.
Before proceeding, let $\alpha^{+}=\inf\{\alpha\in\mathbb{R}:R(\alpha)\geq 0\}$ where $R(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]$. By construction, $\alpha^{+}\geq 0$. By $\mu<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ we also have that $\alpha^{+}<+\infty$. By Lemma C.1, $R$ is a right-continuous function, and therefore $R(\alpha^{+})=\lim_{\alpha\downarrow\alpha^{+}}R(\alpha)\geq 0$. We now claim that $\alpha^{+}>0$. To show this claim, assume $\alpha^{+}=0$. Then, $0\leq R(\alpha^{+})=R(0)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq 0)\mid W_{0}=1]\leq 0$, which implies $\mathbb{P}(\tau_{0}(X)=\mu\mid W_{0}=1)=1$. This is ruled out by the assumption that $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]$. Therefore, $\alpha^{+}>0$.
We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. For all $\underline{w}^{*}$ such that $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\alpha^{+}\mid W_{0}=1]}{\alpha^{+}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{+}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$\leq\frac{\mathbb{E}[1\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[0\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$=\mathbb{E}[\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$\equiv P^{+}.$
The third equality follows from $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$. The inequality follows from $\{0,1\}$ being lower/upper bounds for $\underline{w}^{*}(X)$.
Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{+}$. We now show that this inequality is binding by finding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{+}=1\mid W_{0}=1)=P^{+}$.
Let
$\underline{w}^{+}(X)$ =$=\begin{cases}1&\text{ if }T_{\mu}<\alpha^{+}\\ 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}&\text{ if }T_{\mu}=\alpha^{+}\\ 0&\text{ if }T_{\mu}>\alpha^{+}.\end{cases}$
This function is bounded above by 1 because $R(\alpha^{+})\geq 0$ and $\alpha^{+}>0$. To show it is bounded below by 0, consider cases where $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$ or $R(\alpha^{+})$ equal and differ from 0. If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0$ or $R(\alpha^{+})=0$, then $\underline{w}^{+}(X)\in\{0,1\}\subseteq[0,1]$ and it is therefore bounded below by 0. If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ and $R(\alpha^{+})>0$, then
$1-\frac{R(\alpha^{+})}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$ $=\frac{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
$=\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
$=\frac{-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}.$
By the definition of $\alpha^{+}$ as an infimum, we must have that $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, implying that $R(\alpha)$ is discontinuous at $\alpha^{+}$. By Lemma C.1, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)]$ is left-continuous and satisfies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)]\leq R(\alpha)$. Therefore, since $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, we must have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+}-\varepsilon)]<0$ for all $\varepsilon>0$. Letting $\varepsilon\downarrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})]\leq 0$. Therefore $-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{*}(X)\geq 0$ in that case as well.
Next, we compute
$\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]$ $=\mathbb{E}[\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]$
$\qquad+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{E}[\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]$
$=\mathbb{E}[\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]+\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$=\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$=P^{+}.$
Finally we verify that $\frac{\mathbb{E}[\underline{w}^{+}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]}=\mu$. This condition is equivalent to $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]=0$, which we verify here:
$\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ $=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]$
$\qquad+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{E}[T_{\mu}1(T_{\mu}=\alpha^{+})\mid W_{0}=1]$
$=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$
$=R(\alpha^{+})-R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)$
$=R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0).$
Therefore, $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ equals 0 when $R(\alpha^{+})=0$. When $R(\alpha^{+})>0$, we have that $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ as shown earlier. So $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ is also equal to 0 in this case.
We conclude the proof by showing that $\underline{w}^{+}(X)$ corresponds to $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)$ for some $W^{+}\in\text{SP}(W_{0})$. Let $U\sim\text{Unif}(0,1)$ satisfy $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$ and define
$W^{+}$ $=\left(\mathbbm{1}(T_{\mu}<\alpha^{+})+\mathbbm{1}\left(T_{\mu}=\alpha^{+},U\leq 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\right)\cdot W_{0}.$
By construction, $W^{+}\in\{0,1\}$, $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)=\underline{w}^{+}(X)$, $\mathbb{P}(W^{+}=1\mid W_{0}=0)=0$, and
$W^{+}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$. Also, since $\mu\in\mathcal{S}(\tau_{0};W_{0})$, $\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W_{0}=1)>0$. Since $\alpha^{+}>0$ we have that $\mathbb{P}(W^{+}=1\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}<\alpha^{+}\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)>0$. Therefore $W^{+}$ is a regular subpopulation of $W^{+}$ for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=P^{+}$, hence $P^{+}$ is the maximum.
Case 2: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
As in case 1,
$\overline{P}(a,W_{0};\{\tau_{0}\})$ $\leq\max_{\underline{w}^{*}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
Before proceeding, let $\alpha^{-}=\sup\{\alpha\in\mathbb{R}:L(\alpha)\leq 0\}$ where $L(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]$. By construction, $\alpha^{-}\leq 0$ and by $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ we have that $\alpha^{-}>-\infty$. By Lemma C.1, $L$ is a left-continuous function, and therefore $L(\alpha^{-})=\lim_{\alpha\uparrow\alpha^{-}}L(\alpha)\leq 0$. Similarly to case 1, we can show that $\alpha^{-}<0$.
We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. For all $\underline{w}^{*}$ such that $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\alpha^{-}\mid W_{0}=1]}{\alpha^{-}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{-}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mid W_{0}=1]}{\alpha^{-}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{+}}$
$\leq\frac{\mathbb{E}[1\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{+})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[0\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
$=\mathbb{E}[\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
$\equiv P^{-},$
which follows a similar argument as above. This implies $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{-}$. We now show that this inequality is an equality by finding $W^{-}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{-}=1\mid W_{0}=1)=P^{-}$.
Let
$\underline{w}^{-}(X)$ =$=\begin{cases}1&\text{ if }T_{\mu}>\alpha^{+}\\ 1-\frac{L(\alpha^{-})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)\neq 0)}{\alpha^{-}\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)}&\text{ if }T_{\mu}=\alpha^{-}\\ 0&\text{ if }T_{\mu}<\alpha^{+}.\end{cases}$
The rest of the proof symmetrically follows the one for the previous case.
Case 3: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
Note that $W^{*}=W_{0}$ is the largest regular subpopulation of $W_{0}$. Since $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, we have that $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is trivially maximized at $1$.
Case 4: $\mu\notin\mathcal{S}(\tau_{0};W_{0})$
By Theorem 3.2, there does not exist a regular subpopulation $W^{*}$ satisfying $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ and therefore the supremum equals 0 by its definition. ∎
Proof of Theorem 4.3.
Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Using the law of iterated expectations and letting $\underline{w}^{\prime}(X)=\mathbb{P}(W^{\prime}=1\mid W_{0}=1,X)$, we can write
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\underline{w}^{\prime}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\underline{w}^{\prime}(X)\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)W^{\prime}\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1,W_{0}=1\right]\mathbb{P}(W^{\prime}=1\mid W_{0}=1)}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1,W_{0}=1\right]\mathbb{P}(W^{\prime}=1\mid W_{0}=1)}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
The second equality is valid due to $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$. The third and fourth follow from the law of iterated expectations, and the fifth from $W^{\prime}$ being a subpopulation of $W_{0}$. The second to last line follows from Proposition 3.1 and from $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ implying $\mathbb{P}(\underline{w}^{\prime}(X)>0\mid W_{0}=1)>0$.
Similarly, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu^{\prime}(\underline{w}^{*},\tau_{0})$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1,X)$. Therefore, by Proposition 3.1, we have that $\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right)-\mu^{\prime}(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
Let $\tau^{*}(X)=\frac{a(X)/\underline{w}^{\prime}(X)}{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}$. We have $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$ by construction. Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$ and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
Thus, we must have $\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau^{*}\right)-\mu^{\prime}(\underline{w}^{*},\tau^{*})=0$. Expanding this equality yields
00 0 $=\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau^{*}\right)-\mu^{\prime}(\underline{w}^{*},\tau^{*})$
$=\mathbb{E}\left[\tau^{*}(X)\left(\frac{a(X)/\underline{w}^{\prime}(X)}{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}\right)\mid W^{\prime}=1\right]$
$=\mathbb{E}[\tau^{*}(X)^{2}\mid W^{\prime}=1].$
This implies that $\mathbb{P}(\tau^{*}(X)=0\mid W^{\prime}=1)=1$. In turn, this implies that
$\mathbb{P}\left(a(X)=\frac{\underline{w}^{\prime}(X)\underline{w}^{*}(X)\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}\mid W^{\prime}=1\right)=1.$
We have that $\underline{w}^{*}(X)\geq 0$ and $\underline{w}^{\prime}(X)\geq 0$ almost surely, and $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)>0$ by the assumption that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Also, $\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]=\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1. Therefore, $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$.
[…]
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0},W^{\prime})$, and define
$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}.$
We verify that $W^{*}$ is a regular subpopulation of $W^{\prime}$. First, we see that $\mathbb{P}(W^{*}=1)>0$ because
$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)+\mathbb{P}(W^{*}=1\mid W^{\prime}=0)\mathbb{P}(W^{\prime}=0)$
$=\mathbb{P}\left(\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}=1\mid W^{\prime}=1\right)\mathbb{P}(W^{\prime}=1)$
$=\mathbb{E}\left[\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
$=\mathbb{E}\left[\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
$=\frac{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
$=\frac{\mathbb{E}[a(X)\mid W_{0}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
$>0.$
The second equality follows from $\mathbb{P}(W^{*}=1\mid W^{\prime}=0)=0$, which holds by the construction of $W^{*}$. The fourth equality holds from $U\perp\!\!\!\perp(X,W^{\prime})$ and from $\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\in[0,1]$ almost surely given $W^{\prime}=1$. To establish the final inequality, recall that $\mathbb{P}(W^{\prime}=1)$ and $\mathbb{E}[a(X)\mid W_{0}=1]$ are positive by assumption. Also $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))\leq\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ since $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1)$ is a subset of $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1)$. That $W^{*}$ satisfies the two properties of Definition 3.1 holds immediately. Therefore, $W^{*}$ is a regular subpopulation of $W^{\prime}$.
Finally, let
$\underline{w}^{*}(X)$ $=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$
$=\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)$
$=\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}.$
Letting $\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]=\tau_{0}(X)\in\mathcal{T}_{\text{all}}$ and using Proposition 3.1, we can see that
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{\prime}(X)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
$=\mu(a,\tau_{0}).$
Since $\tau_{0}\in\mathcal{T}_{\text{all}}$ was arbitrary, we have that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ which concludes the proof.
Part 2: $\mathcal{T}=\{\tau_{0}\}$
To simplify the notation in the proof, we let $\mu\equiv\mu(a,\tau_{0})$.
()($\Longrightarrow$)( ⟹ ) First, let $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. Since $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)=0$, which implies $\mathbb{P}(\tau_{0}(X)>\mu\mid W^{\prime}=1)=1$. Since $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$, we can write by Proposition 3.1
$\mu$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
$\geq\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mu\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}=\mu.$ (C.1)
The inequality is strict unless $\mathbb{E}[(\underbrace{\tau_{0}(X)-\mu}_{>0\text{ w.p.1}})\underbrace{\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)}_{\in[0,1]}\mid W^{\prime}=1]=0$ holds. This holds if $\mathbb{P}((\tau_{0}(X)-\mu)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$, which in turns occurs if and only if $\mathbb{P}(\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$. This implies $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)=0$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)=0$, a contradiction of $W^{*}\in\mathcal{W}(a,W^{\prime},\{\tau_{0}\})$. Therefore, the inequality in (C.1) is strict and yields $\mu>\mu$, a contradiction. Therefore, $\mathcal{W}(a;W^{\prime},\{\tau_{0}\})=\emptyset$ when $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$.
[…]
$\mathcal{X}^{-}$ =$=\{x\in\text{supp}(X):\tau_{0}(x)\leq\mu\}\qquad\text{ and }\qquad\mathcal{X}^{+}=\{x\in\text{supp}(X):\tau_{0}(x)\geq\mu\}.$
By $\mu\in\mathcal{S}(\tau_{0};W^{\prime})$, $\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)>0$. Similarly, $\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)>0$.
Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W^{\prime},W_{0})$. For $u\in[0,1]$ let
$W^{*}(u)$ $=(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+}))\cdot W^{\prime}.$
We can see that $W^{*}(u)\in\{0,1\}$, that $W^{*}(u)\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$, and that $\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=0)=0$. To show that $W^{*}(u)$ characterizes a regular subpopulation of $W^{\prime}$, we also show that it is nonzero with positive probability:
$\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)$ $=\mathbb{P}(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+})=1\mid W^{\prime}=1)$
$=\mathbb{P}(U>u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)+\mathbb{P}(U\leq u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)$
$=(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)$
$>0$
for all $u\in[0,1]$. Therefore, $\mathbb{P}(W^{*}(u)=1)=\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)>0$ by $\mathbb{P}(W^{\prime}=1)>0$. Hence, $W^{*}(u)\in\text{SP}(W^{\prime})$ for all $u\in[0,1]$.
For a given $u$, we have that $\underline{w}^{*}(X;u)\equiv\mathbb{P}(W^{*}(u)=1\mid X,W^{\prime}=1)=(1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+})$. Therefore, using Proposition 3.1,
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X;u)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{*}(X;u)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\mid W^{\prime}=1]}$
$=\frac{(1-u)\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]+u\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}.$
By construction, $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{-})\leq\mu\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{+})\geq\mu\mathbbm{1}(X\in\mathcal{X}^{+})$ almost surely. Therefore,
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(0)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}\leq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}=\mu$
and
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(1)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}\geq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}=\mu.$
By the continuity of $\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ in $u$ and the intermediate value theorem, there exists $u^{*}\in[0,1]$ such that $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u^{*})=1]$ and $W^{*}(u^{*})\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. ∎
Proof of Theorem 4.4.
Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
First suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$. From Theorem 4.3, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Written differently, we have that
$\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$ $=\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1\right]}$
for all $\tau_{0}\in\mathcal{T}_{\text{all}}$. From derivations in the proof of Theorem 4.3, we have that $\mathbb{P}(C\cdot\frac{a(X)}{\underline{w}^{\prime}(X)}=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1)=1$ for some positive constant $C>0$.
Since $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\leq 1$ almost surely given $W^{\prime}=1$, we must have $C\cdot a(X)/\underline{w}^{\prime}(X)\leq 1$ almost surely given $W^{\prime}=1$. This means $C$ is bounded above by $\inf(\text{supp}(\underline{w}^{\prime}(X)/a(X)\mid W^{\prime}=1))$, which is strictly positive by assumption. Therefore,
$\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ $=\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$
$\leq\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\right]}{\mathbb{P}(W^{\prime}=1)}$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)$
$\qquad\cdot\mathbb{E}\left[\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid X,W_{0}=1\right]\mid W_{0}=1\right]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\cdot\mathbb{E}[a(X)\mid W_{0}=1]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}.$
The fifth line follows from $W^{\prime}$ being a subpopulation of $W_{0}$, and the last line follows from the law of iterated expectations.
This upper bound is sharp because it is attained by setting
$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}$
and noting that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ from Theorem 4.3.
Now suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$. By Theorem 4.3, $\mathcal{W}(a;W^{\prime},\mathcal{T})=\emptyset$ and therefore $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})$ is zero.
Part 2: $\mathcal{T}=\{\tau_{0}\}$
In this case, we seek to maximize $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ subject to $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. Using Proposition 3.1, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ as $\mathbb{E}[\tau_{0}(X)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]/\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$. The result then follows from a direct application of Theorem 4.2 that replaces $W_{0}$ by $W^{\prime}$ in its statement and $\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ by $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$ in the proofs. ∎
Appendix D Difference-in-Differences
Goodman-Bacon (2021) provides the following representation of the two-way fixed effects estimand under the assumption that group-level average treatment effects are constant over time:
$\beta_{\text{TWFE}}$ $=\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\Bigg{[}\sum_{j=1}^{k-1}\sigma_{jk}^{k}+\sum_{j=k+1}^{K}\sigma_{kj}^{k}\Bigg{]}\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1],$
where
$\sigma_{jk}^{k}\;=\;\frac{\mathbb{P}(G=j)\cdot\mathbb{P}(G=k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}}{\operatorname{var}(D^{\perp\left(G_{t_{1}},\ldots,G_{t_{K-1}},P_{1},\ldots,P_{T}\right)})}$
and
$\sigma_{kj}^{k}\;=\;\frac{\mathbb{P}(G=j)\cdot\mathbb{P}(G=k)\cdot\big{[}1-\mathbb{P}(D=1\mid G=k)\big{]}\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}}{\operatorname{var}(D^{\perp\left(G_{t_{1}},\ldots,G_{t_{K-1}},P_{1},\ldots,P_{T}\right)})}.$
[…]
$\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]}$
$=$ $\frac{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1]}{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)},$
it becomes clear that, for each group $k$ other than the always treated and the never treated,
$a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
$+$ $\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}1-\mathbb{P}(D=1\mid G=k)\big{]}\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]},$
and this, in turn, implies that
$a_{B}(k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$ (D.1)
$+$ $\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}.$
D.1 Equivalence of Weight Functions
We now show that the weights obtained in equation (D.1) are equivalent to those in Proposition 2.5. First, we rewrite the weights in (D.1) as follows:
$a_{B}(k)$ $=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
$+\;\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}$
$=\mathbb{P}(D=1,G<k)-\mathbb{P}(G<k)\mathbb{E}[D\mid G=k]+(1-\mathbb{E}[D\mid G=k])\mathbb{P}(G>k)$
$\quad-\frac{1-\mathbb{E}[D\mid G=k]}{\mathbb{E}[D\mid G=k]}\mathbb{P}(D=1,G>k)$
$=\mathbb{P}(D=1,G<k)-\mathbb{P}(G<k)\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)-\mathbb{E}[D\mid G=k]\mathbb{P}(G>k)$
$\quad-\frac{1}{\mathbb{E}[D\mid G=k]}\mathbb{P}(D=1,G>k)+\mathbb{P}(D=1,G>k)$
$=\mathbb{P}(D=1,G\neq k)-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
$=(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]\mathbb{P}(G=k))-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
$=\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right).$
For $k\in\{2,\ldots,T\}$, the weights in Proposition 2.5 are equal to
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D].$ (D.2)
because they are the average of the weights in Proposition 2.4 conditional on $G=k$. The proof of Proposition 2.5 explicitly shows that
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=\mathbb{P}(D=0\mid G=k)\cdot(\mathbb{P}(D=0\mid P\geq k)+\mathbb{P}(D=1\mid P<g)).$
Let us look at the difference between the weights in (D.1) and (D.2). Fix $k\in\{2,\ldots,T\}$ and write
$\left(1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D]\right)-\left(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)\right)$
$=1-\mathbb{E}[D\mid P\geq k]-\mathbb{P}(G>k)+\frac{\mathbb{E}[D\mathbbm{1}(G>k)]}{\mathbb{E}[D\mid G=k]}$
$=\mathbb{E}[\mathbbm{1}(G\leq k)]-\frac{\mathbb{E}[D\mathbbm{1}(P\geq k)]}{\mathbb{E}[\mathbbm{1}(P\geq k)]}+\frac{\mathbb{E}[D\mathbbm{1}(G>k)]}{\mathbb{E}[\mathbbm{1}(k\leq P)]}$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[D\mathbbm{1}(G>k)]-\mathbb{E}[D\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[\mathbbm{1}(k<G\leq P)]-\mathbb{E}[\mathbb{E}[D\mid P]\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[\mathbb{E}[\mathbbm{1}(k<G\leq P)\mid P]]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[(F_{G}(P)-F_{G}(k))\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]-F_{G}(k)\mathbb{E}[\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
$=0.$
Therefore, the weights in Proposition (2.5), (D.1), and (D.2) are all equal to one another.